### v2

# The chain algebra of a pure poset

Abstract.
We extend the chain algebra, originally defined in [4] for finite distributive lattices to that of finite pure posets. We show this algebra corresponds to the Ehrhart ring of a (0,1)-polytope, termed the chain polytope, and characterize the indecomposability of this polytope. Furthermore, we prove the normality of the chain algebra, describe its canonical module, and extend one of main results from [4] by computing its Krull dimension. For width-2 pure posets, we determine the algebra’s regularity and conditions for it to be Gorenstein or nearly Gorenstein.
Key words and phrases:
Chain algebra, Chain polytope, Canonical module, regularity, Gorenstein, nearly Gorenstein
2010 Mathematics Subject Classification:
Primary 13A02, 05E40 Secondary 13H60.
1. Introduction
The field of studying toric rings defined by combinatorial objects is an active area of research, driven by the dynamic interplay between toric geometry, combinatorics, and commutative algebra. Examples of such algebras include edge rings of graphs [22, 17], base rings of matroids or discrete polymatroids [1, 23, 8, 14], toric rings and Ehrhart rings of lattice polytopes [6], and Hibi rings of posets [12], and rings of stable set polytopes [15].
Recently, a novel type of toric rings, termed the chain algebra of a finite distributive lattice, was introduced and studied in [4]. Let $L$ be a finite distributive lattice and $\mathbb{K}[C_{L}]$ denote its the chain algebra. It was proved in [4] that the toric ideal of $\mathbb{K}[C_{L}]$ has a square-free initial ideal, which implies that it is a normal domain. Furthermore, the Krull dimension of $\mathbb{K}[C_{L}]$ is equal to $|L|-\mathrm{rank}(L)$, where $\mathrm{rank}(L)$ is the size of a maximal chain of $L$ minus one. Additionally, it was shown that the toric ideal of $\mathbb{K}[C_{L}]$ is generated by quadratical binomials if and only if $L$ is plannar. In this case, $\mathbb{K}[C_{L}]$ coincides with the Hibi ring of another lattice derived from $L$. These investigations inspire several questions:
(1) Could we describe the canonical module of such an algebra?
(2) Could we define the concept of chain algebra in a more general setting?
The second question arises from the recognition that the only essential condition in the definition of a chain algebra is that all maximal chains must have the same length, while finite distributive lattices represent a very special type of pure posets. A poset is referred as pure if all its maximal chains have the same length. In this paper, we extend the concept of the chain algebra from finite distributive lattices to pure posets. Let $P$ be a finite pure poset, and let $R:=\mathbb{K}[a|a\in P]$ be the polynomial ring over a field $\mathbb{K}$ with variables corresponding to the elements in $P$. Each maximal chain of $P$ is then associated with a square-free monomial in $R$, which is the product of all the elements in this chain. The chain algebra of $P$, denoted by $\mathbb{K}[C_{P}]$, is defined as the $\mathbb{K}$-subalgebra of $R$ generated by these monomials. To study the properties of $\mathbb{K}[C_{P}]$, we observe that it is the affine semigroup ring of the following affine semigroup.
Definition 1.1.
Let $P$ be a pure poset. We use $C_{P}$ to denote the set of all functions $f$ from $P$ to $\mathbb{N}_{0}:=\{0,1,2,\cdots\}$ such that
$U_{f}:=\prod_{a\in P}a^{f(a)}$
belongs to the chain algebra of $P$.
It is clear that $C_{P}$ is an affine sub-semigroup of $\mathbb{N}_{0}^{P}$ and the chain algebra of $P$ is the affine semigroup ring of $C_{P}$. Here, $\mathbb{N}_{0}^{P}$ denotes the set of all functions from $P$ to $\mathbb{N}_{0}$, which is isomorphic to $\mathbb{N}_{0}^{|P|}$ as semigroups.
Let $D\subset\mathbb{R}^{d}$ be a lattice polytope, which is a convex polytope whose vertices all have integral coordinates. Following [6], the Ehrhart ring $A(D)$ of $D$ is defined by
$A(D):=\mathbb{K}[x_{1}^{a_{1}}\cdots x_{d}^{a_{d}}y^{k}\colon\;(a_{1},\ldots,a_{d})\in kD\cap\mathbb{Z}^{d},k\in\mathbb{N}_{0}].$
Additionally, the toric ring $\mathbb{K}[D]$ of $D$ is defined to be
$\mathbb{K}[D]:=\mathbb{K}[x_{1}^{a_{1}}\cdots x_{d}^{a_{d}}y\colon\;(a_{1},\ldots,a_{d})\in D\cap\mathbb{Z}^{d}].$
If all vectors in $D\cap\mathbb{Z}^{d}$ belongs to a hyperplance, the variable “$y$” in the definition of $\mathbb{K}[D]$ may be removed. It is evident that $A(D)=\mathbb{K}[D]$ if and only if $D$ has the integer decomposition property. Here, a lattice polytope $D$ is said to have the integer decomposition property (IDP) if for all integer $k>0$ and all $X\in kD\cap\mathbb{Z}^{d}$, there exist $Y_{1},\ldots Y_{k}\in D\cap\mathbb{Z}^{d}$ such that $X=Y_{1}+\cdots+Y_{d}$.
To every finite pure poset $P$, we associate a $(0,1)$-polytope, which we call the chain polytope of $P$.
Definition 1.2.
Let $P$ be a finite pure poset. Each maximal chain of $P$ can be associated with a function from $P$ to $\mathbb{N}_{0}$ that maps elements within the chain to 1 and elements outside the chain to 0. Suppose that $f_{1},\ldots,f_{t}\in\mathbb{N}_{0}^{P}$ represents all functions associated to maximal chains of $P$. We define the chain polytope of $P$, denoted by $D_{P}$, as the convex hull of $(0,1)$-vectors $f_{1},\ldots,f_{t}$ contained in $\mathbb{R}^{P}$.
In view of Lemma 1.3, the chain algebra $\mathbb{K}[C_{P}]$ coincides with the toric ring $\mathbb{K}[D_{P}]$.
Lemma 1.3.
Let $f_{1},\ldots,f_{t}$ be $(0,1)$-vectors in $\mathbb{R}^{d}$ such that $|f_{1}|=\cdots=|f_{t}|$. Then
$\mathrm{conv}(f_{1},\ldots,f_{t})\cap\mathbb{Z}^{d}=\{f_{1},\ldots,f_{t}\}.$
Here, for any $f\in\mathbb{R}^{d}$, $|f|$ denotes the number $f(1)+\cdots+f(d).$
Proof.
The case when $|f_{1}|=\cdots=|f_{t}|=2$ was proved in [19], and the approach can be applied in this general case. We include a proof for the completeness. Let $f\in\mathrm{conv}(f_{1},\ldots,f_{t})\cap\mathbb{Z}^{d}$. Then, it is clear that $f$ is also a $(0,1)$-vector. Write $f=a_{1}f_{1}+\cdots+a_{t}f_{t}$, where $a_{1}+\cdots+a_{t}=1$ with $a_{i}\geq 0$ for all $i$. We may assume that $a_{1}>0$ w.l.o.g. We claim that if $f(i)=1$ then $f_{1}(i)=1$ for $i=1,\ldots,d$. In fact, if not, let $j\in[d]$ such that $f(j)=1$ but $f_{1}(j)=0$. Then $f(j)=a_{2}f_{2}(j)+\cdots+a_{t}f_{t}(j)\leq a_{2}+\cdots+a_{t}<1$, a contradiction. This implies we have the inclusion $\{1\leq i\leq d\colon\;f(i)=1\}\subseteq\{1\leq i\leq d\colon\;f_{1}(i)=1\}$. On the other hand, it is clear that $|f|=|\{1\leq i\leq d\colon\;f(i)=1\}|=|\{1\leq i\leq d\colon\;f_{1}(i)=1\}|=|f_{1}|$. Hence, we conclude that $f=f_{1}$, completing the proof.
The paper is organized as follows. Section 2 provides a review of the essential background information on posets, affine semigroups (or rings), canonical modules, nearly Gorenstein, Segre product and other related concepts.
In Section 3, we characterize the elements of $C_{P}$ by using a system of inequalities. As a consequence, we demonstrate that $C_{P}$ is a normal affine semigroup, and thus the chain algebra $\mathbb{K}[C_{P}]$ is a normal domain. In Section 4, based on the characterization of functions of $C_{P}$, we show that the chain polytope $D_{P}$ has the integer decomposition property, and thus the chain algebra $\mathbb{K}[C_{P}]$ is identical to the Ehrhart ring $A(D_{P})$. Additionally, we characterize when $D_{P}$ is an indecomposable polytope in terms of the property of $P$.
In Section 4, we present an explicit description of the canonical module of a chain algebra.
In Section 5, the formula for the Krull dimension of $\mathbb{K}[C_{P}]$ is presented in terms of the combinatorial data of $P$, which generalizes the corresponding result for finite distributive lattices as established in [4]. Note that we employ a very distinct method from [4].
The edge ring of a bipartite graph is identical to the chain algebra of a pure set of rank one. Consequently, certain results pertaining to the edge ring of a bipartite graph can be derived as a specific instance of the above findings.
Nearly Gorensteinness, as a generalization of Gorensteiness, was introduced in [9], and then studied extensively in many papers, such as [6, 10, 13, 16]. Since a chain algebra is normal, and thus Cohen-Macaulay, it is natural to ask when it is Gorenstein or nearly Gorenstein. This question seems difficult in a general setting, even when $P$ is confined to being a distributive lattice. Therefore, we narrow our focus to examining the chain algebra of pure posets with a width of 2.
In the final section, we study the chain algebra of a pure poset of width 2. To accomplish this, we introduce the concept of a basic poset and decompose every pure poset of width 2 into the ordinal sum of basic posets and anti-chains. Based the previously obtained results, we compute the regularity of $\mathbb{K}[C_{P}]$ and provide the conditions under which $\mathbb{K}[C_{P}]$ are Gorenstein or nearly Gorenstein, for pure posets $P$ of width 2.
2. Preliminaries
In this section, we collect concepts and basic facts that are crucial for this paper. We use $\mathbb{K}$ to denote a field, $\mathbb{Z}$ and $\mathbb{Q}$ to represent the ring of integers and the rational field, $\mathbb{N}$ to denote the set of positive integers, and $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$ to represent the set of all natural numbers.
2.1. Marriage Lemma and perfect matching
Let $G$ be a simple graph with vertex set $V(G)$ and edge set $E(G)$. By definition, a perfect matching of $G$ is a subset of $E(G)$ consisting of pairwise disjoint edges such that their union covers all vertices of $G$. A simple graph $G$ is called a bipartite graph (with partition $(V_{1},V_{2})$) if $V(G)$ has a decomposition $V(G)=V_{1}\sqcup V_{2}$ such that for any edge $e\in E(G)$, $|e\cap V_{1}|=|e\cap V_{2}|=1$.
When $G$ is bipartite graph with partition $(V_{1},V_{2})$, we equivalently define a perfect matching of $G$ as a bijection $\phi$ from $V_{1}$ to $V_{2}$ such that each $a\in V_{1}$ is adjacent to its image $\phi(a)\in V_{2}$. For any subset $S\subseteq V(G)$, the neighborhood set $N_{G}(S)$ of $S$ in $G$ is the set of vertices of $G$ that are adjacent to some vertex in $S$. The following lemma, known as the the Marriage Lemma, has several versions. We adapt the version from [7, Lemma 9.1.2].
Lemma 2.1.
Let $G$ be a bipartite graph with partition $(V_{1},V_{2})$. Then $G$ has a perfect matching if and only if $|V_{1}|=|V_{2}|$ and $|S|\leq|N_{G}(S)|$ for all subset $S\subseteq V_{1}$.
2.2. Posets and chain algebras
A partially ordered set (also known as a poset) is a set together with a binary relation $\leq$ that is reflexive, antisymmetric, and transitive. If $x\leq y$ and $x\neq y$, we write $x<y$.
Let $P$ be a finite poset. A chain of $P$ is a subposet of $P$ such that every two elements in it are comparable. The length of a chain $C$ is $\ell(C)=|C|-1$. A chain of $P$ is maximal if it is not strictly contained in any other chain of $P$. Conversely, an anti-chain of $P$ is a subposet of $P$ in which every two distinct elements are incomparable.
Definition 2.2.
A finite poset $P$ is called pure or graded of rank $r$ for some $r\in\mathbb{N}_{0}$ if every maximal chain of $P$ has length $r$.
Let $x,y\in P$ with $x<y$. We say $y$ covers $x$ if there is no $z\in P$ such that $x<z<y$. For a poset $P$, we say that a function $\rho:P\rightarrow\mathbb{N}_{0}$ is a rank function if $\rho$ satisfies the following conditions:
(1) $\rho(m)=0$ for every minimal element $m\in P$, and
(2) $\rho(y)=\rho(x)+1$ for all $x,y\in P$ such that $y$ covers $a$.
Every finite pure poset has an unique rank function. In what follows, we always assume that $P$ is a finite pure poset of rank $r$ with a rank function $\rho$. We use $P_{i}$ to denote the subset of $P$ consisting of elements $a$ such that $\rho(a)=i$ for $i=0,\ldots,r$. For $i=0,\ldots,r-1$, let $G_{i}(P)$ denote the bipartite graph with partition $(P_{i},P_{i+1})$ in which for any $a\in P_{i}$ and $b\in P_{i+1}$, $a$ is adjacent to $b$ if and only if $b$ covers $a$.
Definition 2.3.
Let $P$ be a finite pure poset and let $R$ denote the polynomial ring $\mathbb{K}[P]$. To every maximal chain $C:a_{0}<a_{1}<\cdots<a_{r}$ in $P$, we associate a square-free monomial: $a_{C}=a_{0}a_{1}\cdots a_{r}$ in $R$. The $\mathbb{K}$-subalgebra of $R$ generated by $a_{C}$ for all maximal chains $C$ of $P$ is called the chain algebra of $P$, denoted by $\mathbb{K}[C_{P}]$.
It is clear that $\mathbb{K}[C_{P}]$ is the affine semigroup ring of the affine semigroup $C_{P}$, as defined in Definition 1.1.
Notice that for any $a\in P_{i}$, there is a maximal chain of the form: $a_{0}<a_{1}\cdots<a_{i-1}<a_{i}=a<a_{i+1}<\cdots<a_{r}.$ Consequently, we observe that for any $a\in P$, there is at least one $f\in C_{P}$ such that $f(a)>0$.
2.3. Nearly Gorenstein and pseudo-Gorenstein
Let $R$ be a positively graded $\mathbb{K}$-algebra with maximal graded ideal $\mathfrak{m}$. For an $R$-module $M$, its trace, denoted $\mathrm{tr}_{R}(M)$, is the sum of the ideals $\phi(M)$ with $\phi\in\mathrm{Hom}_{R}(M,R)$. Assume further that $R$ is a Cohen-Macaulay $\mathbb{K}$-algebra admitting a canonical module $w_{R}$. Following [9], $R$ is called nearly Gorenstein if $\mathrm{tr}_{R}(w_{R})\supseteq\mathfrak{m}$. Following [3], $R$ is called pseudo-Gorenstein if $\dim_{\mathbb{K}}(\omega_{R})_{a}=1$, where $a:=\min\{i\colon\;(\omega_{R})_{i}\neq 0\}$. Both nearly Gorenstein and pseudo-Gorenstein properties serve as generalizations of the notion of Gorensteinness.
2.4. Homogenous affine semigroups
Recall from [2] that an affine semigroup $S$ is a finitely generated subsemigroup of $\mathbb{Z}^{d}$ for some $d>0$ with $0\in S$. We only consider affine semigroups that are contained in $\mathbb{N}_{0}^{d}$.
Definition 2.4.
Let $S\subset\mathbb{N}_{0}^{d}$ be an affine semigroup. It is said to be normal if for $\alpha,\beta,\gamma\in S$ and $n>0$ satisfying $n\alpha=n\beta+\gamma$, there is $\gamma^{\prime}\in S$ such that $\gamma=n\gamma^{\prime}$.
A domain is said to be normal if it is integral closed in its fractional field. It is known that $S$ is normal if and only if the toric ring $\mathbb{K}[S]$ is normal, and in this case, $\mathbb{K}[S]$ is Cohen-Macaulay, see [2]. Here, $\mathbb{K}[S]$ is the $\mathbb{K}$-subalgebra of $\mathbb{K}[x_{1},\ldots,x_{q}]$ generated by monomials $X^{\alpha}=x_{1}^{\alpha_{1}}\cdots x_{q}^{\alpha_{q}}$ with $\alpha=(\alpha_{1},\ldots,\alpha_{q})\in S$.
Let $S\subset\mathbb{N}_{0}^{d}$ be a normal affine semigroup. An ideal of $S$ is a subset $T$ of $S$ such that $a+b\in T$ for any $a\in T$ and $b\in S$. Set
$K_{S}=\{\alpha\in S\colon\;\forall\beta\in S,\exists n>0\mbox{ and }\gamma\in S\mbox{ such that }n\alpha=\beta+\gamma\}.$
Then $K_{S}$ is an ideal of $S$. The following lemma is due to [20, (21)].
Lemma 2.5.
Let $S\subset\mathbb{N}_{0}^{d}$ be a normal affine semigroup. Then the canonical module $\omega_{R}$ of the toric ring $R:=\mathbb{K}[S]$ is isomorphic to the ideal $\mathbb{K}[K_{S}]$ of $R$, which by definition is generated by $X^{\alpha}$ with $\alpha\in K_{S}$.
The only affine semigroup we study in this paper is $C_{P}$, where $P$ is a pure poset. As shown in Section 3, $C_{P}$ is a normal affine semigroup. We will use $K_{P}$ to denote $K_{S}$ if $S=C_{P}$.
Recall that an affine semigroup $H$ is homogenous if all its minimal generators lies on an affine hyperplane. In this case, the semigroup ring $\mathbb{K}[H]$ is standard graded by assigning the degree one to all the monomials corresponding to the minimal generators of $H$, and call $\mathbb{K}[H]$ a homogeneous (or standard graded) semigroup ring. Let $H$ be a homogeneous affine semigroups. Then there is a vector $c\in\mathbb{Q}_{0}^{d}$ such that $(c,h)=1$ for any minimal generator $h$ of $H$. If $P$ is a pure poset of rank $r$, it is clear that $C_{P}$ is homogeneous, where $c$ is given by $c(p)=\frac{1}{r+1}$ for all $p\in P$. Set $H_{i}:=\{a\in H\colon\;(c,a)=i\}$ for all $i\geq 0$. Then $H$ can be expressed as the disjoint union $H=\bigsqcup_{i\geq 0}H_{i}$ with $H_{0}=\{\mathbf{0}\}$. Moreover, $H_{i}=H_{1}+\cdots+H_{1}$ for all $i\geq 1$, where $H_{1}$ repeats $i$ times. If $\alpha\in H_{i}$, we write $\deg(\alpha)=i$. It is clear that $\deg(f)=f(P_{i})$ for any $f\in C_{P}$, where $i$ is any number in $[0,r]$.
The $a$-invariant of a standard graded $K$-algebra, denoted by $a(R)$, is defined to be the degree of the Hilbert series of $R$, see [2, Def 4.4.4]. It is known from [5] that if $R$ is Cohen-Macaulay, then $a(R)=-\min\{i\colon\;(\omega_{R})_{i}\neq 0\}$. Hence, if $P$ is a pure poset, then $a(\mathbb{K}[C_{P}])$ is equal to $-\min\{i\colon\;(K_{P})_{i}\neq\emptyset\}.$
2.5. Ordinal sum and Segre product
For $\mathbb{N}_{0}$-graded $\mathbb{K}$-algebras $S$ and $T$, their Segre product $S\#T$ is defined as the $\mathbb{N}_{0}$-graded $K$-algebra: $R:=\bigoplus_{i\geq 0}S_{i}\otimes_{\mathbb{K}}T_{i}.$ If both $T$ and $S$ are Cohen-Macaulay positively graded $K$-algebras admitting canonical modules, then $a(R)=\min\{a(T),a(S)\}$ by [10, Proposition 2.2].
Given disjoint posets $Q_{1}$ and $Q_{2}$, the ordinal sum $Q_{1}\oplus Q_{2}$ of $Q_{1}$ and $Q_{2}$ is defined to be the poset with the underlying set $Q_{1}\sqcup Q_{2}$, where, for any $x_{1},x_{2}\in Q_{1}\sqcup Q_{2}$, $x_{1}\leq x_{2}$ if either $x_{i}\in Q_{i}$ for $i=1,2$ or $\{x_{1},x_{2}\}\subseteq Q_{i}$ for some $i\in\{1,2\}$ and $x_{1}\leq x_{2}$ in the poset $Q_{i}$.
If both $Q_{1}$ and $Q_{2}$ are pure posets, then the ordinal sum $Q_{1}\oplus Q_{2}$ is also a pure poset. Furthermore, its chain algebra $\mathbb{K}[C_{Q_{1}\oplus Q_{2}}]$ is isomorphic to the Segre product $\mathbb{K}[C_{Q_{1}}]\#\mathbb{K}[C_{Q_{2}}]$.
3. Normality
In this section, we describe the functions in $C_{P}$ through a combination of inequalities and equalities. Building upon this description, we show that the chain algebra $\mathbb{K}[C_{P}]$ is a normal domain.
For a finite pure poset $P$ with rank $r$, recall that $P_{i}$ denotes a subset of $P$ consisting of elements of rank $i$ for $0\leq i\leq r$. For a subset $S\subseteq P_{i}$, where $0\leq i<r$, let $N_{P}(S)$ represent the subset of $P_{i+1}$ comprising elements that cover some element in $S$. The main result of this section is as follows:
Theorem 3.1.
Let $P$ be a finite pure poset with rank $r$, and let $f$ be a function from $P$ to $\mathbb{N}_{0}$. Then the following two conditions are equivalent:
  1. $\mathrm{(1)}$
    The function $f$ belongs to the set $C_{P}$.
  2. $\mathrm{(2)}$
    The following two statements hold:
    1. $\mathrm{(a)}$
      For any $0\leq i,j\leq r$,
      $\sum_{a\in P_{i}}f(a)=\sum_{a\in P_{j}}f(a).$
    2. $\mathrm{(b)}$
      For any $10\leq i\leq r-$ 1 and any non-empty subset $V$ of $P_{i}$,
      $\sum_{a\in V}f(a)\leq\sum_{a\in N_{P}(V)}f(a).$
To establish this result, we need some preparations. Let $P$ denote a pure poset, and let $f$ be a function mapping every element of $P$ to $\mathbb{N}_{0}$. We define $S^{f}:=\bigsqcup_{a\in S}\{a^{[1]},\ldots,a^{[f(a)]}\}$ for any subset $S\subseteq P$. By convention, if $f(a)=0$, then $\{a^{[1]},\ldots,a^{[f(a)]}\}$ is understood to be the empty set.
Recall that for $i=0,\ldots,r-1$, $G_{i}(P)$ is the bipartite graph with bipartition $(P_{i},P_{i+1})$, where for $a\in P_{i}$ and $b\in P_{i+1}$, $a$ is adjacent to $b$ if and only if $b$ covers $a$. We now introduce the notation $G_{i}(P)^{f}$ to denote the bipartite graph with bipartition $(P_{i}^{f},P_{i+1}^{f})$, where, for any $a^{[k]}\in P_{i}^{f}$ and any $b^{[\ell]}\in P_{i+1}^{f}$, $a^{[k]}$ is adjacent to $b^{[\ell]}$ if and only if $b$ covers $a$. Following [18, 77.3], we call $G_{i}(P)^{f}$ the $f$-parallelization of $G_{i}(P)$.
Recall that if $G$ is bipartite graph with partition $(V_{1},V_{2})$, a perfect matching of $G$ is defined to be a bijection $\phi$ from $V_{1}$ to $V_{2}$ such that each $a\in V_{1}$ is adjacent to its image $\phi(a)\in V_{2}$.
Lemma 3.2.
Let $f:P\rightarrow\mathbb{N}_{0}$ be a function. Then $U_{f}:=\prod_{a\in P}a^{f(a)}$ belongs to $\mathbb{K}[C_{P}]$ if and only if $G_{i}(P)^{f}$ has a perfect matching for all $i=0,\ldots,r-1$.
Proof.
Define $\Gamma:P^{f}\rightarrow P$ to be the map sending every element of $P^{f}$ to its prototype. In other words, if $a=p^{[j]}$ for some $p\in P$ and for some $1\leq j\leq f(p)$, then $\Gamma(a)=p$. One easily sees that $U_{f}=\prod_{a\in P^{f}}\Gamma(a)$.
Suppose that for $i=0,\ldots,r-1$, the graph $G_{i}(P)^{f}$ has a perfect matching, which we denote by $\phi_{i}$. Write that
$P_{0}^{f}=\{a_{1},a_{2},\ldots,a_{n}\}.$
Let $\varphi_{i}$ denotes the composition $\phi_{i}\phi_{i-1}\cdots\phi_{0}$ for $i=0,\ldots,r-1$. Then
$P_{1}^{f}=\{\varphi_{0}(a_{1}),\varphi_{0}(a_{2}),\ldots,\varphi_{0}(a_{n})\},$
[…]
$P_{i}^{f}=\{\varphi_{i-1}(a_{1}),\varphi_{i-1}(a_{2}),\ldots,\varphi_{i-1}(a_{n})\},$
[…]
$P_{r}^{f}=\{\varphi_{r-1}(a_{1}),\varphi_{r-1}(a_{2}),\ldots,\varphi_{r-1}(a_{n})\}.$
Note that for any $1\leq j\leq n$,
$\Gamma(a_{j})<\Gamma(\varphi_{0}(a_{j}))<\cdots<\Gamma(\varphi_{i}(a_{j}))<\cdots<\Gamma(\varphi_{r-1}(a_{j}))$
is a maximal chain of $P$. Therefore, $U_{f}$ is expressed as the product
$\prod_{j=1}^{n}\Big{(}\Gamma(a_{j})\Gamma(\varphi_{0}(a_{j}))\cdots\Gamma(\varphi_{i}(a_{j}))\cdots\Gamma(\varphi_{r-1}(a_{j}))\Big{)},$
where each term corresponds to a maximal chain of $P$. Consequently, $U_{f}$ belongs to $\mathbb{K}[C_{P}]$.
Conversely, suppose $U_{f}$ belongs to $\mathbb{K}[C_{P}]$. Then there is a positive integer $n\geq 1$ and $n$ maximal chains of $P$, say $a_{0,j}<a_{1,j}<\cdots<a_{r,j}$ for $j=1,\ldots,n$ such that
$U_{f}=\prod_{j=1}^{n}a_{0,j}a_{1,j}\cdots a_{r,j}.$
Denote by $K_{i}$ the multi-set $\{a_{i,1},a_{i,2},\ldots,a_{i,n}\}$ for $i=0,\ldots,r$. Note that
$U_{f}=\prod_{i=0}^{r}\prod_{a\in K_{i}}a=\prod_{i=0}^{r}\prod_{a\in P_{i}^{f}}\Gamma(a).$
Fix $i\in[0,r]$. Since $\prod_{a\in K_{i}}a=\prod_{a\in P_{i}^{f}}\Gamma(a)$, it follows that each $a\in P_{i}$ appears $f(a)$ times in $K_{i}$. Thus, there is a bijection, say $w_{i}$, from $K_{i}$ to $P_{i}^{f}$ such that $\Gamma(w_{i}(a_{i,j}))=a_{i,j}$ for all $j=1,\ldots,n$. For $i=0,\ldots,r-1$, consider the following diagram:
[…]
Here, $b_{i}$ denotes the bijection $K_{i}\rightarrow K_{i+1}$ sending $a_{i,j}$ to $a_{i+1,j}$ for $j=1,\ldots,n$. Since $b_{i},w_{i}$ and $w_{i+1}$ are all bijections, there is a bijection $\phi_{i}:P_{i}^{f}\rightarrow P_{i+1}^{f}$ such that the above diagram is commutative. It remains to show that $\phi_{i}$ is a perfect matching of $G_{i}(P)^{f}$. Let $a\in P_{i}^{f}$. There is an unique $1\leq j\leq n$ such that $a_{i,j}\in K_{i}$ and $w_{i}(a_{i,j})=a$. By the definition of $\phi_{i}$, $\phi_{i}(a)=w_{i+1}(a_{i+1,j}).$ Therefore, $\Gamma(a)=a_{i,j}<a_{i+1,j}=\Gamma(\phi_{i}(a))$, implying $a$ is adjacent in $\phi_{i}(a)$ in $G_{i}(P)^{f}$. Consequently, $\phi_{i}$ is a perfect matching of $G_{i}(P)^{f}$, as required.
To simplify the notation, we set $f(A):=\sum\limits_{a\in A}f(a)$ for all subsets $A\subseteq P$. We now move to prove Theorem 3.1.
Proof.
In this proof, the notation $G_{i}(P)^{f}$ will be written as $G_{i}^{f}$ for short.
(1$\Rightarrow$ (2) Fix $i\in[0,r-1]$. By Lemma 3.2, $G_{i}^{f}$ is a bipartite graph with partition $(P_{i}^{f},P_{i+1}^{f})$ that possesses a perfect matching. From this it follows that $|P_{i}^{f}|=|P_{i+1}^{f}|$ by Lemma 2.1. As a result, we can conclude that $f(P_{i})=f(P_{i+1})$ since $|P_{j}^{f}|=f(P_{j})$ holds for all $j=0,\ldots,r$. This establishes the validity of statement (a).
Now, let $V$ be a non-empty subset of $P_{i}$. We may assume without loss of generality that $f(V)\neq 0$, ensuring that $V^{f}$ is a non-empty subset of $P_{i}^{f}$. In view of Lemma 2.1, we deduce that $|V^{f}|\leq|N_{G_{i}^{f}}(V^{f})|$. Observing that $N_{G_{i}^{f}}(V^{f})\subseteq N_{P}(V)^{f}$ and that $f(N_{P}(V))=|N_{P}(V)^{f}|$, we arrive at the inequality $f(V)\leq f(N_{P}(V))$. Therefore, statement (b) is also proven.
(2$\Rightarrow$ (1): Fix $i\in[0,r-1]$. According to Lemma 3.2, it suffices to demonstrate that $G_{i}^{f}$ possesses a perfect matching under the conditions outlined in (2). Given that $f(P_{i})=f(P_{i+1})$, we deduce that $|P_{i}^{f}|=|P_{i+1}^{f}|$. Consider any nonempty subset $S\subseteq P_{i}^{f}$. Setting $V=\Gamma(S)$, then it holds that
$N_{G_{i}^{f}}(S)=(N_{P}(V))^{f}\quad\text{and}\quad S\subseteq V^{f}.$
By virtue of (2), we have $|V^{f}|=f(V)\leq f(N_{P}(V))$, whence it follows that
$|S|\leq|V^{f}|=f(V)\leq|(N_{P}(V))^{f}|=|N_{G_{i}^{f}}(S)|.$
Consequently, Lemma 2.1 implies that $G_{i}^{f}$ indeed has a perfect matching.
Now, we are ready to prove that the chain algebra $\mathbb{K}[C_{P}]$ is normal.
Corollary 3.3.
Let $P$ be a finite pure poset. Then $C_{P}$ is a normal affine semigroup. Consequently, $\mathbb{K}[C_{P}]$ is a normal domain.
Proof.
Let $f,g,$ and $h$ be functions belonging to $C_{P}$ such that $nf=ng+h$ for some $n>0$. To prove that $C_{P}$ is normal, it suffices to show that $h=nh^{\prime}$ for some $h^{\prime}\in C_{P}$. Define $\alpha=g-f$. For any $x\in P$, we have $\alpha(x)=g(x)-f(x)$ is an integer. Given that $n\alpha(x)=h(x)$ and $h(x)$ is non-negative, it follows that $\alpha(x)\geq 0$. This establishes that $\alpha$ is a function from $P$ to $\mathbb{N}_{0}$. Furthermore, since $\alpha=\frac{h}{n}$ and $h\in C_{P}$, it holds that for all pure subsets $X$ of $P$, $\alpha(X)\leq\alpha(N_{P}(X))$. Additionally, for all $0\leq i,j\leq r$, we have $\alpha(P_{i})=\alpha(P_{j})$. These properties imply $\alpha\in C_{P}$ by Theorem 3.1.
Therefore, since $h=n\alpha$ and $\alpha\in C_{P}$, we conclude that $C_{P}$ is normal. Finally, the normality of $\mathbb{K}[C_{P}]$ follows from the normality of $C_{P}$.
4. Chain polytope of $P$
In this section, we prove the chain polytope of any finite pure poset has the integer decomposition property, and characterize when it is indecomposable.
In view of Subsection 2.4, $C_{P}$ is a homogenous affine semigroup wherein $\deg(f)=f(P_{i})$ for any $i\in[0,r]$. This implies that any $f\in(C_{P})_{k}$ is the sum of $k$ functions from $(C_{P})_{1}$. It is worth noting that $(C_{P})_{1}$ is the set of minimal generators of $C_{P}$.
Proposition 4.1.
Let $P$ be a finite pure poset. Then $D_{P}$ has the integer decomposition property. Consequently, $\mathbb{K}[C_{P}]$ is the same as the Ehrhart ring $A(D_{P})$.
Proof.
Let $f\in kD_{P}\cap\mathbb{Z}^{P}$, where $k$ is any positive integer. To prove that $D_{P}$ has IDP, it suffices to show that $f$ is contained in $(C_{P})_{k}$. We do this in the following steps. First, we notice that that $f$ is a function from $P$ to $\mathbb{N}_{0}$. Next, by Definition 1.2, we may write $f=k(a_{1}f_{1}+\cdots+a_{t}f_{t})$, where $f_{i}\in(C_{P})_{1}$ for all $i=1,\ldots,t$ and the coefficients $a_{i}$ satisfy $a_{1}+\cdots+a_{t}=1$ with $a_{i}\geq 0$ for all $i=1,\ldots,t$. According to Theorem 3.1, for all $i=1,\ldots,t$, we have $f_{i}(P_{j})=1$ for $j=0,\ldots,r$ and $f_{i}(X)\leq f_{i}(N_{P}(X))$ for all pure set $X\subseteq P$. From these, it follows that $f(P_{j})=k$ for $j=0,\ldots,r$ and that
$f(X)=k\sum\limits_{j=1}^{t}a_{j}f_{j}(X)\leq k\sum\limits_{j=1}^{t}a_{j}f_{j}(N_{P}(X))=f(N_{P}(X))$
for all pure set $X\subseteq P$. Thus, $f$ belongs to $(C_{P})_{k}$, as required.
Recall that the direct product of two polytopes $D\subset\mathbb{R}^{d}$ and $E\subset\mathbb{R}^{e}$ is denoted by $D\times E\subset\mathbb{R}^{d+e}$. Let $A,B$ be subsets of $\mathbb{\mathbb{R}}^{d}$. Their Minkowski sum is defined as $A+B:=\{x+y\colon\;x\in A,y\in B\}$.
Following [6], a lattice polytope $D$ is called indecomposable provided that $D$ is not a singleton and any representation of $D$ as a Minkowski sum $D=D_{1}+D_{2}$ of lattice polytopes requires that either $D_{1}$ or $D_{2}$ is a singleton. If $D$ is a (0,1)-polytope, there is an easier approach to check if $D$ is indecomposable due to [6, Lemma 30]: if $D$ is a non-singleton $(0,1)$-polytope, then $D$ is indecomposable precisely when there do not exist non-empty $(0,1)$-polytopes $D_{1}$ and $D_{2}$ such that $D_{1}\times D_{2}$.
Theorem 4.2.
Let $P$ be a finite pure poset with rank $r\geq 1$. Then the following statements are equivalent:
  1. (1)
    The chain polytope $D_{P}$ is indecomposable.
  2. (2)
    $G_{i}(P)$ is not a complete bipartite graph for all $i=0,\ldots,r-1$.
  3. (3)
    If $P$ is the ordinal sum $Q_{1}+Q_{2}$, then either $Q_{1}$ or $Q_{2}$ is empty.
Proof.
The proof of $(2)\Leftrightarrow(3)$ is trivial. If $P$ is the ordinal sum $Q_{1}+Q_{2}$, then $D_{P}=D_{Q_{1}}\times D_{Q_{2}}$. This proves $(3)\Rightarrow(1)$. It remains to prove $(2)\Rightarrow(1)$.
$(2)\Rightarrow(1)$ Suppose on the contrary that $D_{P}$ is indecomposable. Then, according to [6, Lemma 30], there exist $(0,1)$-polytopes $D_{1}\subset\mathbb{R}^{Q_{1}}$ and $D_{2}\subset\mathbb{R}^{Q_{2}}$ such that $D_{P}$ is the direct product $D_{1}\times D_{2}$. Here, $Q_{1}$ and $Q_{2}$ are non-empty subsets of $P$ such that $Q_{1}\cap Q_{2}=\emptyset$ and $Q_{1}\cup Q_{2}=P$.
We use the support of $f$ to refer to the subset $\{x\in P\colon\;f(x)=1\}$ for any $f\in D_{P}\cap\mathbb{Z}^{P}$ or $f\in D_{1}\cap\mathbb{Z}^{Q_{1}}$ or $f\in D_{2}\cap\mathbb{Z}^{Q_{2}}$. It is clear that every $f$ is determined by its support. By $D_{P}=D_{1}\times D_{2}$, it means that if $S$ is the support of a $(0,1)$-vector in $D_{P}$, then $S\cap Q_{1}$ and $S\cap Q_{2}$ are the supports of $(0,1)$-vectors in $D_{1}$ and $D_{2}$ respectively. Conversely, if $S_{1}$ and $S_{2}$ are the supports of $(0,1)$-vectors in $D_{1}$ and $D_{2}$ respectively, then $S_{1}\cup S_{2}$ is the support of a $(0,1)$-vector in $D_{P}$, i.e., a maximal chain of $P$.
We claim that for any $a_{1}\in Q_{1}$ and any $a_{2}\in Q_{2}$, $a_{1}$ and $a_{2}$ are comparable in $P$. Since $P$ is pure, there exist maximal chains $C_{1}$ and $C_{2}$ of $P$ such that $a_{j}\in C_{j}$ for $j=1,2$. Since $Q_{j}\cap C_{1}$ and $Q_{j}\cap C_{2}$ are supports of some (0,1)-vectors of $D_{j}$ for $j=1,2$, the union of $Q_{1}\cap C_{1}$ and $Q_{2}\cap C_{2}$ is a maximal chain of $P$. Given that $a_{i}\in Q_{i}\cap C_{i}$ for $i=1,2$, it follows that $a_{1}$ and $a_{2}$ belongs to a maximal chain of $P$. In particular, $a_{1}$ and $a_{2}$ are comparable in $P$, proving the claim.
By this claim, we have for each $i=0,\ldots,r$, either $P_{i}\subseteq Q_{1}$ or $P_{i}\subseteq Q_{2}$. Since both $Q_{1}$ and $Q_{2}$ are non-empty, there exists $k\in[0,r-1]$ such that $P_{k}\subseteq Q_{1}$ and $P_{k+1}\subseteq Q_{2}$. However, the claim implies that $G_{k}(P)$ is a complete bipartite graph, a contradiction.
A bipartite graph can be regarded as a pure poset of rank one, and its edge algebra is isomorphic to the chain algebra of the corresponding poset.
Example 4.3.
Let $G$ be a connected bipartite graph. By Theorem 4.2, its edge polytope is indecomposable precisely when $G$ is not complete. Hence, according to [6, Theorem 32], the edge ring $\mathbb{K}[G]$ of $G$ is Gorenstein if and only if it is nearly Gorenstein in this case.
5. Canonical modules
In this section we describe the functions from the canonical module of $\mathbb{K}[C_{P}]$. The following definition is essential.
Definition 5.1.
A subset of $P$ is pure if it consists of elements of same rank. For a pure subset $X$ of $P$, we call $X$ to be strict if there is $f\in C_{P}$ such that $f(X)<f(N_{P}(X))$.
Recall that $K_{P}$ is defined as the ideal of $C_{P}$ consisting of all functions $f\in C_{P}$ such that for any $g\in C_{P}$, there exist a positive integer $n$ and an element $h$ in $S$ satisfying the equation $nf=g+h$. By Lemma 5.5, the canonical module of $\mathbb{K}[C_{P}]$ is isomorphic to the ideal of $\mathbb{K}[C_{P}]$ generated by $\prod_{x\in P}x^{f(x)}$ with $f\in K_{P}$. Hence, elucidating the structure of the canonical module of $\mathbb{K}[C_{P}]$ is equivalent to characterizing the functions contained in $K_{P}$.
Lemma 5.2.
Let $f\in C_{P}$. Then $f$ belongs to $K_{P}$ if and only if the following conditions hold:
$\mathrm{(1)}$ $f(X)<f(N_{P}(X))$ for any strict subset $X$ of $P$;
$\mathrm{(2)}$ $f(x)>0$ for all $x\in P$.
Proof.
$\Rightarrow$ Let $f$ be a function belonging to $K_{P}$. Consider any strict subset $X$ of $P$. There exists $g\in C_{P}$ such that $g(X)<g(N_{P}(X))$. Given that $f\in K_{P}$, it follows that there exists $n>0$ and $h\in C_{P}$ such that $nf=g+h$. As a result, we have
$f(N_{P}(X))-f(X)=\frac{1}{n}(g(N_{P}(X))-g(X))+\frac{1}{n}(h(N_{P}(X))-h(X))>0.$
Take an arbitrary element $x\in P$. Since there is at least one maximal chain of $P$ containing $x$, it follows that there is $g^{\prime}\in C_{P}$ such that $g^{\prime}(x)>0$. Since $nf=g^{\prime}+h^{\prime}$ for some $n>0$ and for some $h^{\prime}\in C_{P}$, we obtain $nf(x)=g^{\prime}(x)+h^{\prime}(x)>0$. From this, we conclude that $f(x)>0$.
$\Leftarrow$ Suppose that $f$ is a function in $C_{P}$ that fulfills both conditions (1) and (2). Given an arbitrary function $g\in C_{P}$, it is routine to verify that the function $h:=nf-g$ belongs to $C_{P}$, where $n$ is defined as
$n:=\max\{g(N_{P}(X))-g(X)\colon\;X\mbox{ is a strict set of }P\}+\max\{g(x)\colon\;x\in P\}.$
Therefore, we conclude that $f\in K_{P}$, as desired.
We next identify all the strict subsets of $P$. To this end, we introduce some additional notions.
Fix $i\in\{0,\ldots,r-1\}$. The bipartite graph $G_{i}(P)$ is a disjoint union of its connected components, say $G_{i,1},\ldots,G_{i,k_{i}}$. Note that all $G_{i,j}$ are bipartite graphs and we may assume that $G_{i,j}$ has a bipartition $(P_{i,j},N_{P}(P_{i,j}))$ such that $P_{i,j}\subseteq P_{i}$ for $j=1,\ldots,k_{i}$. One easily notes that $P_{i}=P_{i,1}\sqcup\cdots\sqcup P_{i,k_{i}}$ and $P_{i+1}=N_{P}(P_{i,1})\sqcup\cdots\sqcup N_{P}(P_{i,k_{i}}).$
Definition 5.3.
For all $i=0,\ldots,r-1$ and $j=1,\ldots,k_{i}$, we call $P_{i,j}$ to be a connected component of rank $i$.
It is evident that every connected component of rank $i$ is a pure subset of $P$.
Lemma 5.4.
Let $P$ be a pure poset of rank $r$, and fix an integer $i$ such that $10\leq i\leq r-$ 1.
$\mathrm{(1)}$ Every connected component of rank $i$ is not a strict set.
$\mathrm{(2)}$ Every non-empty proper subset of any connected component of rank $i$ is a strict set.
$\mathrm{(3)}$ Consider any pure subset $X\subseteq P_{i}$. The subset $X$ is a strict set if and only if there exists at least one $j\in\{1,\ldots,k_{i}\}$ such that $X\cap P_{i,j}$ is a strict subset.
Proof.
(1) Fix $f\in C_{P}$. Then $f(P_{i,j})\leq f(N_{P}(P_{i,j}))$ for all $j=1,\ldots,k_{i}$. Furthermore, $f(P_{i})=\sum_{j=1}^{k_{i}}f(P_{i,j})$ and $f(P_{i+1})=\sum_{j=1}^{k_{i}}f(N_{P}({P_{i,j}}))$. Note that $f(P_{i})=f(P_{i+1})$, it follows that $f(P_{i,j})=f(N_{P}(P_{i,j}))$ for all $j=1,\ldots,k_{i}$. Thus, every $P_{i,j}$ is not a strict set.
(2) Let $S$ be a non-empty proper subset of $P_{i,j}$ for some $i,j$. Since the bipartite graph $G_{i,j}$ with bipartition $(P_{i,j},N_{P}(P_{i,j}))$ is connected, there exist $\alpha\in S$ and $\beta\in P_{i,j}\setminus S$ such that $N_{P}(\alpha)\cap N_{P}(\beta)$ is non-empty. Let $\gamma\in N_{P}(\alpha)\cap N_{P}(\beta)$. Then there is at least one maximal chain of the following form:
$\alpha_{0}<\alpha_{1}<\alpha_{2}<\cdots<\alpha_{i-1}<\beta<\gamma<\gamma_{i+1}<\cdots<\gamma_{r}.$
Denote by $w$ the function from $P$ to $\mathbb{N}_{0}$ which corresponds to this chain. That is $w(p)=1$ if $p$ belongs to this chain and $w(p)=0$ otherwise. Since $\gamma$ belongs to $N_{P}(S)$, it follows that
$\sum_{p\in S}w(p)=0\mbox{\qquad and \qquad}\sum_{p\in N_{P}(S)}w(p)\geq w(\gamma)=1.$
It particularly implies $w(S)<w(N_{P}(S))$, and thereby proving $S$ is a strict set.
(3) Set $X_{j}=X\cap P_{i,j}$. Then $X=\sqcup_{j=1}^{k_{i}}X_{j}$ and $N_{P}(X)=\sqcup_{j=1}^{k_{i}}N_{P}(X_{j})$. This implies $f(X)<f(N_{P}(X))$ if and only if there is at least one $j\in[1,k_{i}]$ such that $f(X_{j})<f(N_{P}(X_{j}))$. From this, the result follows.
Combining Lemmas 5.2 and 5.4, we obtain the main result of this section.
Theorem 5.5.
Let $P$ be a pure poset of rank $r$. Then for any $f\in C_{P}$, $f$ belongs to $K_{P}$ if and only if the following conditions are satisfied:
$\mathrm{(1)}$ for any $10\leq i\leq r-$ 1 and for any non-empty proper subset of any connected component of rank $i$, it holds: $f(X)<f(N_{P}(X))$.
$\mathrm{(2)}$ $f(x)>0$ for all $x\in P$.
6. Dimension of a chain algebra
Based on the results obtained in previous sections, in this section we present an explicit formula for the Krull dimension of a chain algebra of any pure poset, which extend the corresponding result in [4]. We begin with a result that may be well-known in linear algebra, and for the sake of the completeness we include a proof.
Lemma 6.1.
Let $V$ be a linear subspace of $\mathbb{Q}^{n}$ and let $S$ be an open subset of $\mathbb{Q}^{n}$. If $S\cap V\neq\emptyset$, then $V$ is spanned by the vectors belonging to $S\cap V$.
Proof.
Let $\alpha_{0}\in S\cap V$ and let $\alpha_{1},\ldots,\alpha_{r}$ be a basis of $V$. Since $\alpha_{0}$ is an interior points of $S$, there is an positive integer $k>0$ such that $\frac{1}{k}(\alpha_{i}-\alpha_{0})+\alpha_{0}$ belongs to $S$ for $i=1,\ldots,r$. Thus, $S\cap V$ contains vectors $\frac{1}{k}(\alpha_{i}-\alpha_{0})+\alpha_{0}$ for $i=1,\ldots,r$ and $\alpha_{0}$. This implies that $\alpha_{1},\ldots,\alpha_{r}$ belong to the space spanned by the vectors belonging to $S\cap V$. Therefore, the result follows.
Theorem 6.2.
Let $P$ be a finite pure poset with rank r. Let $k_{i}$ denote the number of connected components of rank $i$ for $i=0,\ldots,r-1$. Then
$\dim\mathbb{K}[C_{P}]=|P|-(k_{0}+\cdots+k_{r-1}).$
Proof.
For each $p\in P$, let $X_{p}$ denote an indeterminate indexed by $p$. Consider the following systems of equalities and inequalities:
(§) $\sum_{p\in P_{i,j}}X_{p}=\sum_{p\in N_{P}(P_{i,j})}X_{p},\qquad\mbox{ for }i=0,\ldots,r-1,j=1,\ldots,k_{i}.$
(†) $X_{p}\geq 0,\quad\mbox{ for }p\in P.$
(${\dagger}_{<}$) $X_{p}>0,\quad\mbox{ for }p\in P.$
(‡) $\sum_{p\in S}X_{p}\leq\sum_{p\in N_{P}(S)}X_{p},\mbox{ for any proper subset }S\neq\emptyset\mbox{ of }P_{i,j}\mbox{ for some }i,j.$
(${\ddagger}_{<}$) $\sum_{p\in S}X_{p}<\sum_{p\in N_{P}(S)}X_{p},\mbox{ for any proper subset }S\neq\emptyset\mbox{ of }P_{i,j}\mbox{ for some }i,j.$
It is natural to identify every rational or integral solution to part or all of these equalities or inequalities with a function mapping from $P$ to the set of rational numbers or integral numbers. Under this identification, we have
[…]
  • $K_{P}$ consists of all integral solutions that satisfy $(\S),({\dagger}_{<})$ and $({\ddagger}_{<})$.
For a subset $M$ of the $\mathbb{Q}$-subspace $\mathbb{Q}^{P}$ (the set of all functions from $P$ to $\mathbb{Q}$), we denote by $\langle M\rangle_{\mathbb{Q}}$ the $\mathbb{Q}$-subspace of $\mathbb{Q}^{P}$ generated by $M$. It is evident that the Krull dimension of $\mathbb{K}[C_{P}]$ is equal to $\dim_{\mathbb{Q}}\langle C_{P}\rangle_{\mathbb{Q}}$. We also introduce the following notions:
  • Let $V$ denote the set of all rational solutions to (§)§($\S$)( § );
  • Let $U$ denote the set of all rational solutions to $({\dagger}_{<})$ and $({\ddagger}_{<})$;
[…]
It is straightforward to observe that $V\cap U\subseteq A\subseteq V$. Since $K_{P}\subseteq V\cap U$, it follows that $V\cap U$ is non-empty. Note that $V$ is a $\mathbb{Q}$-subspace and $U$ is an open set of $\mathbb{Q}^{P}$. Therefore, according to Lemma 6.1, $\langle V\cap U\rangle_{\mathbb{Q}}$ coincides with $V$. Given that $\langle C_{P}\rangle_{\mathbb{Q}}=\langle A\rangle_{\mathbb{Q}}$, it follows that the Krull dimension of $\mathbb{K}[C_{P}]$ is equal to the dimension $\dim_{\mathbb{Q}}V$. Finally, observing that the rows of the coefficient matrix of homogeneous system of linear equations (§)§($\S$)( § ) are linear independent (in fact, the coefficient matrix of homogeneous system of linear equations (§)§($\S$)( § ) is in row echelon form), the result follows from the basic linear algebra.
It was proved in [4, Theorem 3.4] that if $L$ is a finite distributive lattice of rank $r$, then the chain algebra $\mathbb{K}[C_{L}]$ has the dimension $|L|-r$. Since every $G_{i}(P)$ is connected when $P$ is a finite distributive lattice, this result is a special case of Theorem 6.2.
7. Posets of width 2
Recall the width of a poset is the maximal size of an anti-chain contained in this poset. Let $P$ be a pure poset of width 2. In this section, we will compute the regularity of $\mathbb{K}[C_{P}]$ and characterize the conditions under which $\mathbb{K}[C_{P}]$ is Gorenstein or nearly Gorenstein.
Hereafter, we always assume that $P$ is a pure poset of width 2 and rank $r\geq 1$ unless otherwise stated. As previously, we denote by $P_{i}$ the set of elements of $P$ with rank $i$, and by $G_{i}(P)$ the bipartite graph with bipartition $(P_{i},P_{i+1})$, where for $a\in P_{i}$ and $b\in P_{i+1}$, $a$ is adjacent to $b$ if and only if $a<b$. It is clear that $P_{i}$ contains at most 2 elements for each $i$. We make the following observations:
(1) If, for some $i$, $G_{i}(P)$ is not connected, then $G_{i}(P)$ consists of two disjoint edges. We may assume $P_{i}=\{a_{1},a_{2}\}$ and $P_{i+1}=\{b_{1},b_{2}\}$ with $a_{j}<b_{j}$ for $j=1,2$. In this case, if we identify $a_{j}$ with the $b_{j}$ for $j=1,2$ to obtain a smaller poset $P^{\prime}$, then it is evident that $C_{P}$ is isomorphic to $C_{P^{\prime}}$, and $K_{P}$ is isomorphic to $K_{P^{\prime}}$. In particular, we have $\mathbb{K}[C_{P}]$ is isomorphic to $\mathbb{K}[C_{P^{\prime}}]$. Hereafter, we denote by $\overline{P}$ the poset that is obtained from $P$ by a series of such identifications, where all corresponding bipartite graphs are connected. Roughly speaking, $\overline{P}$ is obtained from $P$ by removing the disconnected $G_{i}(P)^{\prime}s$.
(2) If there is an index $0<i<r$ such that $G_{i}(P)$ is complete bipartite, then $P$ is the ordinal sum of the subposets $Q_{1}:=P_{0}\cup\cdots\cup P_{i}$ and $Q_{2}:=P_{i+1}\cup\cdots\cup P_{r}$. Furthermore, if $|P_{i}|=1$, then $\mathbb{K}[C_{Q_{1}}]$ is isomorphic to $\mathbb{K}[C_{Q_{1}^{\prime}}]$, where $Q_{1}^{\prime}$ denotes the subposet $Q_{1}\setminus P_{i}$, and similarly, if $|P_{i+1}|=1$, then $\mathbb{K}[C_{Q_{2}}]$ is isomorphic to $\mathbb{K}[C_{Q_{2}^{\prime}}]$, where $Q_{2}^{\prime}:=Q_{2}\setminus P_{i+1}$.
The following definition is now natural.
Definition 7.1.
A pure poset $P$ is called basic if it is of wideth 2 and rank $\geq 1$, and every $G_{i}(P)$ is a connected bipartite graph that is not complete.
The above discussions establish that, given any pure poset $P$ with width 2, we have a poset $\overline{P}$ such that $G_{i}(\overline{P})$ is connected for all possible $i$ and $\mathbb{K}[C_{P}]$ is isomorphic to $\mathbb{K}[C_{\overline{P}}]$. Furthermore, $\overline{P}$ is the ordinal sum of basic posets and posets of rank 00 and width 2. Note that a poset of rank 00 and width 2 is nothing but an anti-chain containing two elements.
In what follows, whenever $P$ is a basic poset, we always assume that $P=\{X_{0},\ldots,X_{r},Y_{0},\ldots,Y_{r}\}$ such that $X_{0}<X_{1}<\cdots<X_{r}$ and $Y_{0}<\cdots<Y_{r}$ are two maximal chains of $P$. For $X\in P$, we say $X$ has degree $i$ if $|N_{P}(X)|=i$, meaning that there are exactly $i$ elements that covers $X$. By our assumption, for each $i\in[0,r-1]$, exactly one of $\{X_{i},Y_{i}\}$ has degree one. Furthermore, if $X_{i}$ has degree one, then $N_{P}(X_{i})=\{X_{i+1}\}$ and $N_{P}(Y_{i})=\{X_{i+1},Y_{i+1}\}$, and similarly, if $Y_{i}$ has degree one then $N_{P}(Y_{i})=\{Y_{i+1}\}$ and $N_{P}(X_{i})=\{X_{i+1},Y_{i+1}\}$. Thus, To fully describe a basic poset, it is enough to identify all the elements that have degree one. We associate every basic poset a sequence of positive integers as follows.
Definition 7.2.
Let $(c_{1},c_{2},\ldots,c_{n})$ be an integral vector, where $c_{1}+\cdots+c_{n}=r$ with $c_{i}\geq 1$ for all $i=1,\ldots,n$. Set $b_{0}=0$ and $b_{i}=c_{1}+c_{2}+\cdots+c_{i}$ for $i=1,\ldots,n$. Let $P$ be a basic poset. By relabelling if necessary, we may assume $X_{0}$ has degree one. We say a basic poset $P$ belongs to type $(c_{1},c_{2},\ldots,c_{n})$ if the degree-one elements of $P$ are as follows:
$X_{0},\ldots,X_{c_{1}-1},Y_{c_{1}},\ldots,Y_{c_{1}+c_{2}-1},\ldots,$
$X_{b_{2i}},X_{b_{2i}+1},\ldots,X_{b_{2i+1}-1},Y_{b_{2i+1}},Y_{b_{2i+1}+1},\ldots,Y_{b_{2i+2}-1},\ldots$
$W_{b_{n-1}},W_{b_{n-1}+1},\ldots,W_{b_{n}-1}.$
Here, $W=X$ if $n$ is odd, and $W=Y$ if $n$ is even.
Up to isomorphisms, every basic poset is determined by its type. Specializations of Theorems 3.1 and 5.5 to basic posets produces the following result.
Lemma 7.3.
Let $P$ be a basic poset of type $(c_{1},\ldots,c_{n})$ with $c_{0}$ set to 0. Consider a function $f:P\rightarrow\mathbb{N}_{0}$ such that $f(P_{i})=f(P_{j})$ for all $0\leq i,j\leq r$. Then
[…]
  • If $i$ is even, then $f(X_{k})\leq f(X_{k+1})$;
  • If $i$ is odd, then $f(X_{k})\geq f(X_{k+1})$.
[…]
$f(X_{0})\leq\cdots\leq f(X_{c_{1}})\geq f(X_{c_{1}+1})\geq\cdots\geq f(X_{c_{1}+c_{2}})\leq\ldots\geq\cdots f(X_{r-1})?f(X_{r}).$
Here, ???? is $\leq$ if $n$ is odd and ???? is $\geq$ if $n$ is even.
[…]
  • If $i$ is even, then $f(X_{k})<f(X_{k+1})$;
  • If $i$ is odd, then $f(X_{k})>f(X_{k+1})$.
[…]
$f(X_{0})<\cdots<f(X_{c_{1}})>f(X_{c_{1}+1})>\cdots>f(X_{c_{1}+c_{2}})<\ldots>\ldots f(X_{r-1})?f(X_{r}).$
Here, ???? is $<$ if $n$ is odd and ???? is $>$ if $n$ is even.
Henceforth, we shall use $\lambda(P)$ (or simply $\lambda$) to represent the maximum value among $c_{i}$ for $1\leq i\leq n$. Additionally, we set $b_{0}=0$, and $b_{i}=c_{1}+\ldots+c_{i}$ for $i=1,\ldots,r$. Recall that for a function $f\in C_{P}$, the degree of $f$, denoted by $\deg(f)$, is defined as $f(P_{i})$, where $i$ can be any number within $[0,r]$. The following lemma allows us to construct functions in $K_{P}$ satisfying certain properties conveniently.
Lemma 7.4.
[…]
Proof.
We extend the definition of $f$ in the following way: For $0\leq i<s$ or $t\leq i\leq n$,
  • If $i$ is even, we define $f(X_{b_{i}})=1,f(X_{b_{i}+1})=2,\ldots,f(X_{b_{i+1}-1})=c_{i}$.
  • If $i$ is odd, we define $f(X_{b_{i}})=\lambda+1$ when $i\neq t$. It is worth noting that if $i=t$, we do not need to define $f(X_{b_{i}})$. Furthermore, we specify the values of $f(X_{b_{i}+1}),\ldots,f(X_{b_{i+1}-1})$ in such a way that they form a strictly decreasing sequence within the interval $[2,\lambda]$. As an example, such a sequence could be $\lambda,\lambda-1,\ldots,\lambda-c_{i}+2$.
  • For all $0\leq i\leq r$, we define $f(Y_{i})=d-f(X_{i})$.
Note that if either $c_{i+1}=1$ or $i=n$, we do not need to define $f(X_{b_{i}+1}),\ldots,f(X_{b_{i+1}-1})$.
By applying Lemma 7.3, it can be easily verified that $f$ indeed belongs to $K_{P}$ and has a degree of $d$, thereby completing the proof.
[…]
Lemma 7.5.
Let $P$ be a basic poset of type $(c_{1},\ldots,c_{n})$. Then $a(\mathbb{K}[C_{P}])=-\lambda-2$.
Proof.
We first show that $\deg(f)\geq\lambda+2$ for all $f\in K_{P}$. Let $f\in K_{P}$. There exists $1\leq i\leq n$ such that $c_{i}=\lambda$. We may assume that $i$ is even w.l.o.g. Set $b_{0}=0$ and $b_{j}=c_{1}+\cdots+c_{j}$ for $j=1,\ldots,n$. Then
$1\leq f(X_{b_{i}})<\cdots<f(X_{b_{i}+c_{i}}).$
This implies $f(X_{b_{i}+c_{i}})\geq\lambda+1$ and so $\deg(f)=f(X_{b_{i}+c_{i}})+f(Y_{b_{i}+c_{i}})\geq\lambda+2$.
By Lemma 7.4, there is a function in $K_{P}$, say $g$, such that $g(X_{b_{i}+j})=1+j$ for $j=0,\ldots,c_{i}=\lambda$ with degree $\lambda+2$. This completes the proof.
Based on this result, we can compute the regularity of the chain algebra of all pure posets of width 2. Recall that the Hilbert series of the chain algebra $\mathbb{K}[C_{P}]$ can be expressed as a rational function $\frac{Q(t)}{(1-t)^{d}}$, where $d$ is equal to the Krull dimension of $\mathbb{K}[C_{P}]$. Since $\mathbb{K}[C_{P}]$ is Cohen-Macaulay, the regularity of $\mathbb{K}[C_{P}]$ is equal to the degree of $Q(t)$. By definition, the invariant $a(\mathbb{K}[C_{P}])$ is the degree of its Hilbert series, which is given by $\deg(Q(t))-d$. Consequently, we have the equation:
$\mathrm{reg}(\mathbb{K}[C_{P}])=d+a(\mathbb{K}[C_{P}]).$
The dimension $d$ can be determined using Theorem 6.2. As discussed in the previous section, the $a$-invariant of the chain algebra of the ordinal sum of two posets $P_{1}$ and $P_{2}$ is given by $a(\mathbb{K}[C_{P_{1}\oplus P_{2}}])=\min\{a(\mathbb{K}[C_{P_{1}}]),a(\mathbb{K}[C_{P_{2}}])\}.$ It is known that every pure poset of width 2 can be expressed as the ordinal sum of basic posets and posets of rank 0 and width 2. If $P$ is a poset of rank 0 and width 2, it is straightforward to see that $\mathbb{K}[C_{P}]$ is Gorenstein with an $a$-invariant of $-2$. Consequently, if $P$ is a pure poset of width 2, the $a$-invariant of $\mathbb{K}[C_{P}]$ can be determined using these observations along with Lemma 7.5. The above discussions provide an approach to compute the regularity of the chain algebra of any pure poset of width 2.
Finally, we will characterize when $\mathbb{K}[C_{P}]$ is Gorenstein or nearly Gorenstein whenever $P$ is a pure poset of width 2. Note that for any $f\in C_{P}$, $\deg(f)$ is equal to $f(P_{i})$ for any $i\in[0,r]$.
Lemma 7.6.
Let $P$ be a simple of type $(c_{1},\ldots,c_{n})$. Then the following statements are equivalent:
  1. (1)
    $\mathbb{K}[C_{P}]$ is pseudo-Gorenstein.
  2. (2)
    There exists an unique function belonging to $K_{P}$ with degree $\lambda+2$.
  3. (3)
    It holds that $c_{i}\in\{1,\lambda\}$ for all $i=1,\ldots,n$, and that if $c_{i}=1$ then $1<i<n$ and $c_{i-1}=c_{i+1}=\lambda$.
Proof.
$(1)\Leftrightarrow(2)$ By Lemma 7.5, we have $a(\mathbb{K}[C_{P}])=-\lambda-2$. Due to this fact, the equivalence of (1) and (2) follows by the definition of pseudo-Gorensteinness.
$(3)\Rightarrow(2)$ Let $f$ be an arbitrary function in $K_{P}$ with $\deg(f)=\lambda+2$. To establish the result, it suffices to prove that the values of $f(X)$ for all $X\in P$ are uniquely determined by the sequence $c_{1},\ldots,c_{n}$. We proceed through the following steps:
Step 1: Given that $c_{1}=\lambda$, we invoke Lemma 7.3 to conclude:
$11\leq f(X_{0})<f(X_{1})<\cdots<f(X_{\lambda})\leq\lambda+$ 1 .
Consequently, it must be that $f(X_{j})=j+1$ for $j=0,\ldots,\lambda$.
Step 2: Consider now the case where $c_{i+1}=\lambda$ for some $i$. Utilizing a similar argument as in Step 1, we deduce:
  • If $i$ is even, then $f(X_{b_{i}+j})=j+1$ for $j=0,\ldots,\lambda$.
  • If $i$ is odd, then $f(X_{b_{i}+j})=\lambda+1-j$ for $j=0,\ldots,\lambda$.
In particular, for each $i$ such that $c_{i+1}=\lambda$, the values of $f$ on the subset
$\{X_{b_{i}},X_{b_{i}+1},\ldots,X_{b_{i}+c_{i+1}-1},X_{b_{i+1}}\}$
of $P$ are uniquely specified.
Step 3: For indices $i$ where $c_{i+1}=1$, note that $c_{i}=c_{i+2}=\lambda$. By the results of Step 2, the values of $f$ at $X_{b_{i}}$ and $X_{b_{i+1}}$ are already uniquely determined. Therefore, the values of $f(X_{i})$ for $i=0,\ldots,r$ are uniquely determined. Since $f(X_{i})+f(Y_{i})=\lambda+2$, it follows that the value of $f(Y_{i})$ is also uniquely determined. In conclusion, the values of $f(X)$ for all $X\in P$ are uniquely specified, completing the proof.
$(2)\Rightarrow(3)$ Suppose on the contrary that there is $1\leq i\leq n$ such that $1<c_{i}<\lambda$. If $i$ is odd, we define functions $f$ and $g$ on the subset $\{X_{b_{i-1}},X_{b_{i-1}+1}\ldots,X_{b_{i-1}+c_{i}-1},X_{b_{i}}\}$ of $P$ by $f(X_{b_{i+1}+j})=j+1$ for $j=0,1,\ldots,c_{i}-1$ and $f(b_{i})=\lambda+1$, and $g(X_{b_{i+1}})=1$, $g(X_{b_{i+1}+j})=j+2$ for $j=1,\ldots,c_{i}-1$ and $g(b_{i})=\lambda+1$. By Lemma 7.4, both $f$ and $g$ could be extended to functions belonging to $K_{P}$ with degree $\lambda+2$. This is a contradiction, implying that if $i$ is odd, $c_{i}$ must be either 1 or $\lambda$. Analogously, we have $c_{i}$ must be either 1 or $\lambda$ if $i$ is even. Hence, we conclude that $c_{i}\in\{1,\lambda\}$ for all $i=1,\ldots,n$.
Suppose that $c_{1}=1$. Then, in view of Lemma 7.4, the functions $f$ and $g$ on the subset$\{X_{0},X_{1}\}$ defined by $f(X_{0})=1,f(X_{1})=\lambda+1$ and $g(X_{0})=2,g(X_{1})=\lambda+1$ can be extended two distinct functions in $K_{P}$ with a degree of $\lambda+2$. This is a contradiction, thereby proving that $c_{1}=\lambda$. Similarly, we can show that $c_{n}=1$.
It remains to show that there is no $22\leq i\leq n-$ 2 such that $c_{i}=c_{i+1}=1$. Suppose on the contrary that $c_{i}=c_{i+1}=1$ for some $22\leq i\leq n-$ 2. If $i$ is odd, according to Lemma 7.4, there are functions $f$ and $g$ of degree $\lambda+2$ such that $f(X_{b_{i-1}})=1,f(X_{b_{i}})=2$ and $f(X_{b_{i+1}})=1$, and that $g(X_{b_{i-1}})=1,g(X_{b_{i}})=\lambda+1$ and $g(X_{b_{i+1}})=1$. This is impossible. Similarly, the case when $i$ is even is also impossible.
Proposition 7.7.
Let $P$ be a basic poset of type $(c_{1},\ldots,c_{n})$. Then the following statements are equivalent:
$\mathrm{(1)}$ the chain algebra $\mathbb{K}[C_{P}]$ is nearly Gorenstein.
$\mathrm{(2)}$ the chain algebra $\mathbb{K}[C_{P}]$ is Gorenstein.
$\mathrm{(3)}$ it holds that $c_{1}=\cdots=c_{n}=\lambda.$
Proof.
$(1)\Leftrightarrow(2)$ By Theorem 4.2, $D_{P}$ is indecomposable, and by Proposition 4.1, $\mathbb{K}[C_{P}]$ coincides with the Ehrhart ring of $D_{P}$. Now, the result follows from [6, Theorem 32].
(2) $\Rightarrow$ (3) Suppose on the contrary that $c_{1}=\cdots=c_{n}=\lambda$ does not hold. By lemma‘7.6, there is $1<k<n$ such that $c_{k}=1$ and $c_{k-1}=c_{k+1}=\lambda\geq 2$. We may assume $k$ is even without loss of generality. By Lemma 7.4 there is a function $\gamma\in K_{P}$ of degree $12\lambda+$ 1 satisfying
$\gamma(X_{b_{k-2}})=1,\gamma(X_{b_{k-2}+1})=2,\cdots,\gamma(X_{b_{k-1}})=\lambda+1,\gamma(X_{b_{k}})=\lambda,$
and
$\gamma(X_{b_{k}+1})=\lambda+1,\gamma(X_{b_{k}+1})=\lambda+2,\ldots,\gamma(X_{b_{k+1}})=2\lambda.$
Let $f$ be the unique function in $K_{P}$ with degree $\lambda+2$. Then $f(X_{b_{k-1}})=\lambda+1,f(X_{b_{k}})=1$. Note that $(\Gamma-f)(X_{b_{k-1}})=0$ and $(\Gamma-f)(X_{b_{k}})=\lambda-1$, we have $\gamma-f$ does not belong to $C_{P}$, a contradiction.
(3) $\Rightarrow$ (2) Let $f$ be the unique function in $K_{P}$ with $\deg(f)=\lambda+2$. Then the sequence $f(X_{0}),f(X_{1}),\ldots,f(X_{r})$ is as follows:
$1,2,\ldots,\lambda,\lambda+1,\lambda,\lambda-1,\ldots,1,2\ldots.$
From this sequence, we observe that $f(X_{i})-f(X_{i+1})\in\{1,-1\}$ for all $i=0,\ldots,r-1$. To prove $\mathbb{K}[C_{P}]$ is Gorenstein, it suffices to show that for any $g\in K_{P}$, the function $g-f$ belongs to $C_{P}$. Take an arbitrary function $g$ from $K_{P}$. According to Lemma 7.3, $f(X_{i})>f(X_{i+1})$ if and only if $g(X_{i})>g(X_{i+1})$, and $f(X_{i})<f(X_{i+1})$ if and only if $g(X_{i})<g(X_{i+1})$ for all $i=0,\ldots,r-1$. Since $g(X_{i})>0$ for all $i=0,\ldots,r$, it follows that $g(X_{i})\geq f(X_{i})$ for all $i$, and so $h:=g-f$ is a function from $P$ to $\mathbb{N}_{0}$.
Now, if $g(X_{i+1})>g(X_{i})$, then $g(X_{i+1})-g(X_{i})\geq f(X_{i+1})-f(X_{i})=1$, and so $h(X_{i+1})\geq h(X_{i})$. Similarly, if $g(X_{i+1})<g(X_{i})$, then $g(X_{i})-g(X_{i+1})\geq f(X_{i})-f(X_{i+1})=1$, and so $h(X_{i+1})\leq h(X_{i})$. Therefore, $h$ and $g$ share the same monotonicity, and so $h$ belongs to $C_{P}$ by Lemma 7.3, as desired.
We are now ready to prove the last main result of the paper. Recalling [10, Theorem 2.7], we have if $R$ is the Segre product $R_{1}\#R_{2}\#\cdots\#R_{n}$ of Gorenstein standard graded $K$-algebras with $\mathbf{a}(R_{i})=-a_{i}$ for $i=1,\ldots,n$, then $R$ is Gorenstein if and only if $a_{1}=\cdots=a_{n}$. Furthermore, $R$ is nearly Gorenstein if and only if $|a_{i}-a_{j}|\leq 1$ for all $1\leq i,j\leq n$.
Example 7.8.
Let $P$ be a pure poset of rank $r$ for which $G_{i}(P)$ is a complete bipartite graph for all $i=0,\ldots,r-1$. Then $P$ is the ordinal sum of the subposets $P_{0},P_{1},\ldots,$ and $P_{r}$. For each $i=0,\ldots,r$, since $P_{i}$ is an anti-chain, $\mathbb{K}[C_{P_{i}}]$ is a polynomial ring with $|P_{i}|$ variables. In particular, $\mathbb{K}[C_{P_{i}}]$ is Gorenstein with $a(\mathbb{K}[C_{P_{i}}])=-|P_{i}|$. Therefore, according to [10, Theorem 2.7], we have $C_{P}$ is Gorenstein if and only if $|P_{0}|=\cdots=|P_{r}|$. Additionally, $C_{P}$ is nearly Gorenstein if and only if there is an integer $d\geq 1$ such that $|P_{i}|$ is equal to either $d$ or $d+1$ for each $i=0,\ldots,r$.
Theorem 7.9.
Let $P$ be a pure poset with a width of 2, and denote by $\overline{P}$ the poset derived from $P$ by eliminating the disconnected bipartite graphs $G_{i}(P)$. Suppose that $\overline{P}$ can be expressed as the ordinal sum of basic posets $P_{1},\ldots,P_{s}$ and posets $Q_{1},\ldots,Q_{t}$ with rank 0 and width 2. Furthermore, assume that each $P_{i}$ for $i=1,\ldots,s$ belongs to type $(c_{i,1},\ldots,c_{i,n_{i}})$.
$\mathrm{(1)}$ If $s=0$, then $\mathbb{K}[P]$ is always Gorenstein.
$\mathrm{(2)}$ If $t>0$ and $s>0$, then $\mathbb{K}[P]$ is not Gorenstein. Moreover, $\mathbb{K}[P]$ is nearly Gorenstein if and only if $c_{i,j}=1$ for all $i=1,\ldots,s,j=1,\ldots,n_{i}$.
$\mathrm{(3)}$ If $t=0$, then the following statements hold:
$\mathrm{(a)}$ $\mathbb{K}[P]$ is Gorenstein if and only if there is a positive integer $\lambda$ such that $c_{i,j}=\lambda$ for all $i=1,\ldots,s,j=1,\ldots,n_{i}$.
$\mathrm{(b)}$ $\mathbb{K}[P]$ is nearly Gorenstein if and only if there is a positive integer $\lambda$ and a subset $A\subseteq[1,s]$ such that $c_{i,j}=\lambda$ for $i\in A$ and $j=1,\ldots,n_{i}$, while $c_{i,j}=\lambda+1$ for $i\in[1,s]\setminus A$ and $j=1,\ldots,n_{i}$.
Proof.
Note that for each $i$, $\mathbb{K}[C_{Q_{i}}]$ is Gorenstein with an a-invariant equal to -2. Therefore, these conclusions are drawn by combining Propositions 7.7, and [10, Theorem 2.7].
We remark that conclusion (1) of the above theorem is actually a special case of Example 7.8.
Acknowledgment: This project is supported by NSFC (No. 11971338). We would like to express our sincere gratitude to Dr. Sora Miyashita, whose comments on the initial version of this paper brought us to the reference [6], resulting in the introduction and the exploration of the chain polytope in this revised version.