### v3

# The chain algebra of a pure poset

Abstract.
We extend the chain algebra, originally defined in [4] for finite distributive lattices to that of finite pure posets. We show this algebra is the Ehrhart ring of a (0,1)-polytope, termed the chain polytope, and characterize the indecomposability of this polytope. Furthermore, we prove the normality of the chain algebra, describe its canonical module, and extend one of main results from [4] by computing its Krull dimension. For width-2 pure posets, we determine the algebra’s regularity and conditions for it to be Gorenstein or nearly Gorenstein.
Key words and phrases:
chain algebra, chain polytope, chain semigroup, pure poset, Canonical module, regularity, Gorenstein, nearly Gorenstein
2010 Mathematics Subject Classification:
Primary 13A02, 05E40 Secondary 13H60.
1. Introduction
The field dedicated to studying toric rings defined by combinatorial objects is a vibrant area of research, fueled by the dynamic interplay among toric geometry, combinatorics, and commutative algebra. Illustrative examples of such algebras include edge rings of graphs [25, 18], base rings of matroids or discrete polymatroids [1, 26, 8, 15], toric rings and Ehrhart rings of lattice polytopes [6], Hibi rings of posets [13], rings of stable set polytopes [16], and cycle algebras of matroids [21, 22].
Recently, a novel type of toric rings, termed the chain algebra of a finite distributive lattice, was introduced and studied in [4]. Let $L$ be a finite distributive lattice and $\mathbb{K}[C_{L}]$ denote its the chain algebra. It was proved in [4] that the toric ideal of $\mathbb{K}[C_{L}]$ has a square-free initial ideal, which implies that it is a normal domain. Furthermore, the Krull dimension of $\mathbb{K}[C_{L}]$ is equal to $|L|-\mathrm{rank}(L)$, where $\mathrm{rank}(L)$ is the size of a maximal chain of $L$ minus one. Additionally, it was shown that the toric ideal of $\mathbb{K}[C_{L}]$ is generated by quadratical binomials if and only if $L$ is plannar. In such a situation, $\mathbb{K}[C_{L}]$ coincides with the Hibi ring of another lattice derived from $L$. These investigations and results inspire the following questions:
(1) Could we describe the canonical module of a chain algebra?
(2) Could we relate a chain algebra to a polytope, given that polytopes are a powerful tool to study algebraic structures?
(3) Could we define the concept of chain algebra in a more general setting?
The third question arises from the recognition that the only essential condition in the definition of a chain algebra is that all maximal chains must have the same length, while finite distributive lattices represent a very special type of pure posets. A poset is called pure if all its maximal chains have the same length. In this paper, we extend the concept of the chain algebra from finite distributive lattices to pure posets. The formal definitions are given below.
Definition 1.1.
Let $P$ be a finite pure poset, and let $R:=\mathbb{K}[x|x\in P]$ be the polynomial ring over a field $\mathbb{K}$. Each maximal chain of $P$ is then associated with a square-free monomial in $R$, which is the product of all the elements in this chain. The chain algebra of $P$, denoted by $\mathbb{K}[C_{P}]$, is defined as the $\mathbb{K}$-subalgebra of $R$ generated by the square-free monomials associated with the maximal chains of $P$. Additionally, we use $C_{P}$ to denote the set of all functions $f$ from $P$ to $\mathbb{N}_{0}:=\{0,1,2,\cdots\}$ such that the monomial $U_{f}:=\prod_{x\in P}x^{f(x)}$ is an element of the chain algebra $\mathbb{K}[C_{P}]$. We call $C_{P}$ the chain semigroup of $P$, since it is an affine sub-semigroup of $\mathbb{N}_{0}^{P}$.
Here and thereafter, for sets $A$ and $B$, the notation $A^{B}$ denotes the set of all maps from $B$ to $A$. Consequently, $\mathbb{N}_{0}^{P}$, equipped with a natural addition operation, is isomorphic to $\mathbb{N}_{0}^{|P|}$, as semigroups. It is clear that the chain algebra $\mathbb{K}[C_{P}]$ is the affine semigroup ring of $C_{P}$. The objective of this paper is to investigate the algebraic properties of the chain algebra and chain semigroup, with a particular focus on their relationship to the combinatorial characteristics of the poset $P$. In our exploration, polytopes emerge as a valuable tool.
Definition 1.2.
Let $P$ be a finite pure poset. For each maximal chain $C$ of $P$, the characteristic function $\chi_{C}$ is the function from $P$ to $\mathbb{R}$ such that $\chi_{C}(x)=1$ if $x\in C$ and $\chi_{C}(x)=0$ otherwise. We then define chain polytope of $P$, denoted by $D_{P}$, as the convex hull in $\mathbb{R}^{P}$ of all these characteristic functions $\chi_{C}$.
Let $D\subset\mathbb{R}^{d}$ be a lattice polytope, which is a convex polytope whose vertices all have integral coordinates. Following [6], the Ehrhart ring $A(D)$ of $D$ is defined by
$A(D):=\mathbb{K}[x_{1}^{a_{1}}\cdots x_{d}^{a_{d}}y^{k}\colon\;(a_{1},\ldots,a_{d})\in kD\cap\mathbb{Z}^{d},k\in\mathbb{N}_{0}].$
Additionally, the toric ring $\mathbb{K}[D]$ of $D$ is defined to be
$\mathbb{K}[D]:=\mathbb{K}[x_{1}^{a_{1}}\cdots x_{d}^{a_{d}}y\colon\;(a_{1},\ldots,a_{d})\in D\cap\mathbb{Z}^{d}].$
They are both $\mathbb{K}$-subalgebras of the polynomial ring $\mathbb{K}[x_{1},\ldots,x_{d},y]$ over a field $\mathbb{K}$. If all vectors in $D\cap\mathbb{Z}^{d}$ belongs to a hyperplance, the variable “$y$” in the definition of $\mathbb{K}[D]$ may be removed. In general, one has $\mathbb{K}[D]\subseteq A(D)$, and it is not difficult to see that $A(D)=\mathbb{K}[D]$ if and only if $D$ has the integer decomposition property. Here, a lattice polytope $D$ is said to have the integer decomposition property (IDP) if for all integer $k>0$ and all $\alpha\in kD\cap\mathbb{Z}^{d}$, there exist $\beta_{1},\ldots\beta_{k}\in D\cap\mathbb{Z}^{d}$ such that $\alpha=\beta_{1}+\cdots+\beta_{k}$. In view of Lemma 2.3, the chain algebra $\mathbb{K}[C_{P}]$ coincides with the toric ring $\mathbb{K}[D_{P}]$.
The first concern of this paper is to depict the functions within $C_{P}$. To achieve this, we associate a pure poset $P$ with a sequence of bipartite graphs $G_{i}(P)$ for $i$ ranging from 0 to $r-1$, where $r$ denotes the length of the longest chain in $P$. Analogous to parallelization in graph theory, we create $f$-parallelizations of these graphs for every function $f$ from $P$ to $\mathbb{N}_{0}$, yielding a new series of bipartite graphs. We establish that $f$ belongs to $C_{P}$ if and only if each graph in its $f$-parallelization possesses a perfect matching. Based on this result, we then use the Marriage Lemma to derive a system of linear inequalities that characterize the functions in $C_{P}$. For details, refer to Theorem 3.1.
Theorem 3.1 yields two immediate implications. Firstly, it asserts that $C_{P}$ is a normal affine semigroup, which implies the chain algebra $\mathbb{K}[C_{P}]$ is a normal domain. We remark that, for finite distributive lattices $L$, the normality of $\mathbb{K}[C_{L}]$ has been established in [4], through a different way: by showing the defining ideal of $\mathbb{K}[C_{L}]$ has a square-free initial ideal. Secondly, it follows that the chain polytope $D_{P}$ has the integer decomposition property, indicating the chain algebra $\mathbb{K}[C_{P}]$ is identical to the Ehrhart ring $A(D_{P})$.
The concept of indecomposable lattice polytope was introduced in [6]. A remarkable property of such polytopes is presented in [6, Theorem 32], which states if $D$ is an indecomposable lattice polytope, then the Enrhart ring $A(D)$ is Gorenstein if and only if it is nearly Gorenstein. We show in Theorem 4.2 that $D_{P}$ is indecomposable if and only if $P$ can not be expressed as the ordinal sum of two non-empty sub-posets of $P$.
Since the chain algebra $\mathbb{K}[C_{P}]$ is a normal domain, according to [2, Theorem 6.3.5], the canonical module $\omega_{\mathbb{K}[C_{P}]}$ of $\mathbb{K}[C_{P}]$ is isomorphic to an ideal of $\mathbb{K}[C_{P}]$, which is itself generated by an ideal of $C_{P}$. We denote this ideal by $K_{P}$. By invoking a formula given in Stanley [23], we can characterize the functions within $K_{P}$ through a system of linear inequalities.
Surprisingly, the characterization of functions within $K_{P}$ is crucial for computing the (Krull) dimension of a chain algebra. Notably, the dimension of the chain algebra $\mathbb{K}[C_{P}]$ is equal to the dimension of the $\mathbb{Q}$-linear space spanned by the functions in $C_{P}$ by [9, Proposition 3.1]. We show that the this space coincides with the one spanned by the functions in $K_{P}$. Recognizing this shared space as the solution space of a homogeneous linear system with a row staircase coefficient matrix, a neat formula for the the dimension of a chain algebra is obtained. For details, see Theorem 6.2. This result generalizes the corresponding result for finite distributive lattices as established in [4]. As a consequence, a formula for the dimension of $D_{P}$ is derived.
The edge ring of a bipartite graph can be regarded as the chain algebra of a pure poset of rank one. Consequently, certain results pertaining to the edge ring of a bipartite graph can be derived as a specific instance of the above findings.
To delve deeply into the algebraic properties of chain algebras, such as regularity, Gorensteinness, and near-Gorensteinness, we must narrow our focus. A significant, yet simpler, class of pure posets to consider is those of width 2. Recall that the width of a poset is defined as the maximum size of its anti-chains. In the final section, we examine the chain algebra associated with a pure poset of this type. To do so, we introduce the concept of basic posets and decompose any pure poset of width 2 into an ordinal sum of basic posets and anti-chains. Leveraging results from previous sections, we compute the regularity of $\mathbb{K}[C_{P}]$ and establish the conditions for $\mathbb{K}[C_{P}]$ to be Gorenstein or nearly Gorenstein when $P$ is a pure poset of width 2. These findings indicate that studying these properties in the general case is a highly challenging task.
The paper is organized as follows. Section 2 reviews background information. Section 3 characterizes the functions from $C_{P}$ and proves that $\mathbb{K}[C_{P}]$ is a normal domain. Section 4 shows $D_{P}$ has the integer decomposition property and characterizes when $D_{P}$ is an indecomposable polytope. Section 5 describes the canonical module of $\mathbb{K}[C_{P}]$. Section 6 presents the Krull dimension formula for $\mathbb{K}[C_{P}]$. The concluding section focuses on chain algebras of pure posets of width 2.
2. Preliminaries
We use $\mathbb{K}$ to denote a field, $\mathbb{Z}$ and $\mathbb{Q}$ to represent the ring of integers and the rational field, $\mathbb{N}$ to denote the set of positive integers, and $\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$ to represent the set of all natural numbers. In this section, we collect concepts and basic facts that are crucial for this paper.
2.1. Marriage Lemma and perfect matching
Let $G$ be a simple graph with vertex set $V(G)$ and edge set $E(G)$. By definition, a perfect matching of $G$ is a subset of $E(G)$ consisting of pairwise disjoint edges such that their union covers all vertices of $G$.
A simple graph $G$ is called a bipartite graph (with partition $(V_{1},V_{2})$) if $V(G)$ has a decomposition $V(G)=V_{1}\sqcup V_{2}$ such that for any edge $e\in E(G)$, $|e\cap V_{1}|=|e\cap V_{2}|=1$. When $G$ is a bipartite graph with partition $(V_{1},V_{2})$, we equivalently define a perfect matching of $G$ as a bijection $\phi$ from $V_{1}$ to $V_{2}$ such that each $a\in V_{1}$ is adjacent to its image $\phi(a)\in V_{2}$. For a subset $S\subseteq V(G)$, the neighborhood set $N_{G}(S)$ of $S$ in $G$ is the set of vertices of $G$ that are adjacent to some vertex from $S$. The following lemma, known as the the Marriage Lemma, has several versions. We adapt the version from [7, Lemma 9.1.2].
Lemma 2.1.
Let $G$ be a bipartite graph with partition $(V_{1},V_{2})$. Then $G$ has a perfect matching if and only if $|V_{1}|=|V_{2}|$ and $|S|\leq|N_{G}(S)|$ for all subset $S\subseteq V_{1}$.
2.2. Posets and chain algebras
A partially ordered set (also known as a poset) is a set together with a binary relation $\leq$ that is reflexive, antisymmetric, and transitive. If $x\leq y$ and $x\neq y$, we write $x<y$.
Let $P$ be a finite poset. A chain of $P$ is a subposet of $P$ such that every two elements in it are comparable. The length of a chain $C$ is $\ell(C)=|C|-1$. A chain of $P$ is maximal if it is not strictly contained in any other chain of $P$. Conversely, an anti-chain of $P$ is a subposet of $P$ in which every two distinct elements are incomparable.
Definition 2.2.
A finite poset $P$ is called pure or graded of rank $r$ for some $r\in\mathbb{N}_{0}$ if every maximal chain of $P$ has length $r$.
Let $x,y\in P$ with $x<y$. We say $y$ covers $x$ if there is no $z\in P$ such that $x<z<y$. For a poset $P$, a function $\rho:P\rightarrow\mathbb{N}_{0}$ is called a rank function if $\rho$ satisfies the following conditions:
(1) $\rho(m)=0$ for every minimal element $m\in P$, and
(2) $\rho(y)=\rho(x)+1$ for all $x,y\in P$ such that $y$ covers $a$.
Every finite pure poset has an unique rank function. In what follows, we always assume that $P$ is a finite pure poset of rank $r$ with a rank function $\rho$. For $0\leq i\leq r$, we denote by $P_{i}$ the subset of $P$ consisting of elements $a$ with rank $i$, i.e., $\rho(a)=i$. For $i=0,\ldots,r-1$, we define $G_{i}(P)$ as the bipartite graph with partition $(P_{i},P_{i+1})$ in which for any $a\in P_{i}$ and $b\in P_{i+1}$, $a$ is adjacent to $b$ if and only if $b$ covers $a$.
Given a semigroup (monoid) $M$, we say that $M$ is generated by a subset $A\subseteq M$ if every element of $M$ is an $\mathbb{N}_{0}$-linear combination of elements of $A$. Any minimal generating set of $M$ is called a Hilbert basis of $M$. Let $P$ be a finite pure poset and let $R$ denote the polynomial ring $\mathbb{K}[x|x\in P]$. To every maximal chain $C:x_{0}<x_{1}<\cdots<x_{r}$ in $P$, we associate a square-free monomial: $x_{C}:=x_{0}x_{1}\cdots x_{r}$ in $R$. By Definition 1.1, the chain algebra $\mathbb{K}[C_{P}]$ is the $\mathbb{K}$-subalgebra of $R$ generated by $x_{C}$ for all the maximal chains $C$ of $P$, and the chain semigroup $C_{P}$ has a unique Hilbert basis that consists of characteristic functions of the maximal chains of $P$.
Notice that for any $x\in P_{i}$, there is a maximal chain of the form: $x_{0}<x_{1}\cdots<x_{i-1}<x_{i}=x<x_{i+1}<\cdots<x_{r}.$ Consequently, we observe that for any $x\in P$, there is at least one $f\in C_{P}$ such that $f(x)>0$.
2.3. Lattice Polytopes
A polytope $P\subset\mathbb{R}^{d}$ is the convex hull of finite vectors of $\mathbb{R}^{d}$. The dimension of a polytope is defined as the dimension of its affine hull. A lattice polytope is a polytope in which all the vertices have integer coordinates and a (0,1)-polytope is a special type of lattice polytope where all the vertices have coordinates that are either 0 or 1. The chain polytope $D_{P}$ is a (0,1)-polytope.
The following result is utilized in numerous literatures, including [22], without explicit mention. As we are unable to trace its original source, we state it here explicitly and give a short proof.
Lemma 2.3.
Let $f_{1},\ldots,f_{t}$ be $(0,1)$-vectors in $\mathbb{R}^{d}$ such that $|f_{1}|=\cdots=|f_{t}|$. Then
$\mathrm{conv}(f_{1},\ldots,f_{t})\cap\mathbb{Z}^{d}=\{f_{1},\ldots,f_{t}\}.$
Here, $\mathrm{conv}(f_{1},\ldots,f_{t})$ denotes the convex hull of $f_{1},\ldots,f_{t}$, and for any $f\in\mathbb{R}^{d}$, $|f|$ denotes the number $f(1)+\cdots+f(d).$
Proof.
The case when $|f_{1}|=\cdots=|f_{t}|=2$ was proved in [20], and the approach can be applied in this general case. We include a proof for completeness. Let $f\in\mathrm{conv}(f_{1},\ldots,f_{t})\cap\mathbb{Z}^{d}$. We may write $f=a_{1}f_{1}+\cdots+a_{t}f_{t}$, where $a_{1}+\cdots+a_{t}=1$ with $a_{i}\geq 0$ for all $i$. From this expression, it follows that $0\leq f(i)\leq 1$ for all $1\leq i\leq d$ and so $f$ is a $(0,1)$-vector. Since $a_{1}+\cdots+a_{t}=1$, there exists at least one $1\leq i\leq t$ such that $a_{i}>0$. Without loss of generality, we assume that $a_{1}>0$. We now claim that if $f(i)=1$ then $f_{1}(i)=1$ for $i=1,\ldots,d$. In fact, if not, there is $1\leq j\leq d$ such that $f(j)=1$ but $f_{1}(j)=0$. This implies $11=f(j)=a_{2}f_{2}(j)+\cdots+a_{t}f_{t}(j)\leq a_{2}+\cdots+a_{t}<$ 1. This contradiction establishes the claim. Consequently, we have the inclusion $\{1\leq i\leq d\colon\;f(i)=1\}\subseteq\{1\leq i\leq d\colon\;f_{1}(i)=1\}$. On the other hand, it is clear that $|f|=|\{1\leq i\leq d\colon\;f(i)=1\}|=|\{1\leq i\leq d\colon\;f_{1}(i)=1\}|=|f_{1}|$. Hence, we conclude that $f=f_{1}$, completing the proof.
According to [9, Definition 4.3], a lattice polytope $D$ is said to have the integer decomposition property (IDP) if for all integer $k>0$ and all $\alpha\in kD\cap\mathbb{Z}^{d}$, there exist $\beta_{1},\ldots\beta_{k}\in D\cap\mathbb{Z}^{d}$ such that $\alpha=\beta_{1}+\cdots+\beta_{k}$. Let $A,B$ be subsets of $\mathbb{\mathbb{R}}^{d}$. Their Minkowski sum is defined as $A+B:=\{x+y\colon\;x\in A,y\in B\}$. Following [6], a lattice polytope $D$ is called indecomposable provided that $D$ is not a singleton and any representation of $D$ as a Minkowski sum $D=D_{1}+D_{2}$ of lattice polytopes requires that either $D_{1}$ or $D_{2}$ is a singleton. Recall that the direct product of two polytopes $D\subset\mathbb{R}^{d}$ and $E\subset\mathbb{R}^{e}$ is defined by $D\times E:=\{(a,b)\colon\;a\in D,b\in E\}\subset\mathbb{R}^{d+e}$. In the case when $D$ is a (0,1)-polytope, [6, Lemma 30] provides an equivalent condition for indecomposability: if $D$ is a non-singleton $(0,1)$-polytope, then $D$ is indecomposable precisely when there do not exist non-empty $(0,1)$-polytopes $D_{1}$ and $D_{2}$ such that $D_{1}\times D_{2}$.
2.4. Nearly Gorenstein and pseudo-Gorenstein
Let $R$ be a positively graded $\mathbb{K}$-algebra with maximal graded ideal $\mathfrak{m}$. For an $R$-module $M$, its trace, denoted $\mathrm{tr}_{R}(M)$, is the sum of the ideals $\phi(M)$ with $\phi\in\mathrm{Hom}_{R}(M,R)$. Assume further that $R$ is a Cohen-Macaulay $\mathbb{K}$-algebra admitting a canonical module $w_{R}$. Following [10], $R$ is called nearly Gorenstein if $\mathrm{tr}_{R}(w_{R})\supseteq\mathfrak{m}$. Following [3], $R$ is called pseudo-Gorenstein if $\dim_{\mathbb{K}}(\omega_{R})_{a}=1$, where $a:=\min\{i\colon\;(\omega_{R})_{i}\neq 0\}$. Both nearly Gorenstein and pseudo-Gorenstein properties serve as generalizations of the notion of Gorensteinness. Nearly Gorensteinness was studied extensively in many papers, such as [6, 11, 14, 17].
2.5. Homogenous affine semigroups
Recall from [2] that an affine semigroup $S$ is a finitely generated sub-semigroup of $\mathbb{Z}^{d}$ for some $d>0$ with $\mathbf{0}\in S$. We only consider positive affine semigroups that are contained in $\mathbb{N}_{0}^{d}$.
Definition 2.4.
Let $S\subset\mathbb{N}_{0}^{d}$ be an affine semigroup. It is said to be normal if for $\alpha,\beta,\gamma\in S$ and $n>0$ satisfying $n\alpha=n\beta+\gamma$, there is $\gamma^{\prime}\in S$ such that $\gamma=n\gamma^{\prime}$.
A domain is said to be normal if it is integral closed in its fractional field. For an affine semigroup $S\subseteq\mathbb{N}_{0}^{d}$, the (affine) semigroup ring $\mathbb{K}[S]$ is defined to be the $\mathbb{K}$-subalgebra of the polynomial ring $\mathbb{K}[x_{1},\ldots,x_{d}]$ generated by monomials $X^{\alpha}:=x_{1}^{\alpha_{1}}\cdots x_{d}^{\alpha_{q}}$ with $\alpha=(\alpha_{1},\ldots,\alpha_{d})\in S$. It is known from [2] that $\mathbb{K}[S]$ is a normal domain if and only if $S$ is a normal affine semigroup, and in this case, $\mathbb{K}[S]$ is Cohen-Macaulay.
Let $S\subset\mathbb{N}_{0}^{d}$ be a normal affine semigroup. An ideal of $S$ is a subset $T$ of $S$ such that $a+b\in T$ for any $a\in T$ and $b\in S$. Set
$K_{S}:=\{\alpha\in S\colon\;\forall\beta\in S,\exists n>0\mbox{ and }\gamma\in S\mbox{ such that }n\alpha=\beta+\gamma\}.$
Then $K_{S}$ is an ideal of $S$. The following lemma is due to [23, (21)].
Lemma 2.5.
Let $S\subset\mathbb{N}_{0}^{d}$ be a normal affine semigroup. Then the canonical module $\omega_{R}$ of the semigroup ring $R:=\mathbb{K}[S]$ is isomorphic to the ideal $\mathbb{K}[K_{S}]$ of $R$, which by definition is generated by $X^{\alpha}$ with $\alpha\in K_{S}$.
The only affine semigroup we study in this paper is $C_{P}$, where $P$ is a pure poset. As shown in Section 3, $C_{P}$ is a normal affine semigroup. We will use $K_{P}$ to denote $K_{S}$ if $S=C_{P}$.
Recall that an affine semigroup $H$ is homogenous if all its minimal generators lies in an affine hyperplane. In this case, the semigroup ring $\mathbb{K}[H]$ is standard graded by assigning the degree one to all the monomials corresponding to the minimal generators of $H$, and call $\mathbb{K}[H]$ a homogeneous (or standard graded) semigroup ring. Let $H$ be a homogeneous affine semigroups. Then there is a vector $c\in\mathbb{Q}^{d}$ such that $(c,h)=1$ for any minimal generator $h$ of $H$. If $P$ is a pure poset of rank $r$, it is clear that $C_{P}$ is homogeneous, where $c$ is given by $c(p)=\frac{1}{r+1}$ for all $p\in P$. Set $H_{i}:=\{a\in H\colon\;(c,a)=i\}$ for all $i\geq 0$. Then $H$ can be expressed as the disjoint union $H=\bigsqcup_{i\geq 0}H_{i}$ with $H_{0}=\{\mathbf{0}\}$. Moreover, $H_{i}=H_{1}+\cdots+H_{1}$ for all $i\geq 1$, where $H_{1}$ repeats $i$ times. When $\alpha\in H_{i}$, we say that $\alpha$ has degree $i$ and write it as $\deg(\alpha)=i$. Returning to chain semigroups, $C_{P}$ is a homogenous affine semigroup with $(C_{P})_{1}$ serving as the unique Hilbert basis of $C_{P}$.
The $a$-invariant of a standard graded $\mathbb{K}$-algebra $R$, denoted by $a(R)$, is defined to be the degree of the Hilbert series of $R$, see [2, Def 4.4.4]. It is known from [5] that if $R$ is a Cohen-Macaulay algebra admitting a canonical module $\omega_{R}$, then $a(R)=-\min\{i\colon\;(\omega_{R})_{i}\neq 0\}$. Hence, if $P$ is a pure poset, then $a(\mathbb{K}[C_{P}])$ is equal to $-\min\{i\colon\;(K_{P})_{i}\neq\emptyset\}.$
2.6. Ordinal sum and Segre product
For $\mathbb{N}_{0}$-graded $\mathbb{K}$-algebras $S$ and $T$, their Segre product $S\#T$ is defined as the $\mathbb{N}_{0}$-graded $\mathbb{K}$-algebra: $R:=\bigoplus_{i\geq 0}S_{i}\otimes_{\mathbb{K}}T_{i}.$ If both $T$ and $S$ are Cohen-Macaulay positively graded $\mathbb{K}$-algebras admitting canonical modules, then $a(R)=\min\{a(T),a(S)\}$ by [11, Proposition 2.2].
Given disjoint posets $Q_{1}$ and $Q_{2}$, the ordinal sum $Q_{1}\oplus Q_{2}$ of $Q_{1}$ and $Q_{2}$ is defined to be the poset with the underlying set $Q_{1}\sqcup Q_{2}$, where, for any $x_{1},x_{2}\in Q_{1}\sqcup Q_{2}$, $x_{1}\leq x_{2}$ if either $x_{i}\in Q_{i}$ for $i=1,2$, or $\{x_{1},x_{2}\}\subseteq Q_{i}$ for some $i\in\{1,2\}$ and $x_{1}\leq x_{2}$ in the poset $Q_{i}$.
If both $Q_{1}$ and $Q_{2}$ are pure posets, then the ordinal sum $Q_{1}\oplus Q_{2}$ is also a pure poset. Furthermore, the chain algebra $\mathbb{K}[C_{Q_{1}\oplus Q_{2}}]$ is isomorphic to the Segre product $\mathbb{K}[C_{Q_{1}}]\#\mathbb{K}[C_{Q_{2}}]$.
3. Chain monoid $C_{P}$ and its Normality
In this section, we describe the functions in $C_{P}$ through a combination of inequalities and equalities. Building upon this description, we show that the chain algebra $\mathbb{K}[C_{P}]$ is a normal domain.
For a finite pure poset $P$ with rank $r$, recall that $P_{i}$ denotes the subset of $P$ consisting of elements of rank $i$ for $0\leq i\leq r$. For a subset $S\subseteq P_{i}$, where $10\leq i<r-$ 1, let $N_{P}(S)$ represent the subset of $P_{i+1}$ comprising elements that cover some element in $S$. Given a function $f\in\mathbb{N}_{0}^{P}$ and a subset $\emptyset\neq A\subseteq P$, for simplicity, we set $f(A):=\sum_{x\in A}f(x)$. The main result of this section is as follows.
Theorem 3.1.
Let $P$ be a finite pure poset with rank $r$, and let $f$ be a function from $P$ to $\mathbb{N}_{0}$. Then the following two conditions are equivalent:
  1. $\mathrm{(1)}$
    The function $f$ belongs to the chain semigroup $C_{P}$.
  2. $\mathrm{(2)}$
    The following two statements hold:
    1. $\mathrm{(a)}$
      For any $0\leq i,j\leq r$,
      $f(P_{i})=f(P_{j}),\mbox{ i.e. }\sum_{x\in P_{i}}f(x)=\sum_{x\in P_{j}}f(x).$
    2. $\mathrm{(b)}$
      For any $10\leq i\leq r-$ 1 and any non-empty subset $V$ of $P_{i}$,
      $f(V)\leq f(N_{P}(V)),\mbox{ i.e. }\sum_{x\in V}f(x)\leq\sum_{x\in N_{P}(V)}f(x).$
To establish this result, we need some preparations. Let $P$ denote a pure poset, and let $f$ be a function mapping every element of $P$ to $\mathbb{N}_{0}$. We define $S^{f}:=\bigsqcup_{x\in S}\{x^{[1]},\ldots,x^{[f(x)]}\}$ for any subset $S\subseteq P$. By convention, if $f(x)=0$, then $\{x^{[1]},\ldots,x^{[f(x)]}\}$ is understood to be the empty set.
Recall that for $i=0,\ldots,r-1$, $G_{i}(P)$ is the bipartite graph with bipartition $(P_{i},P_{i+1})$, where for $a\in P_{i}$ and $b\in P_{i+1}$, $a$ is adjacent to $b$ if and only if $b$ covers $a$. We now introduce the notation $G_{i}(P)^{f}$ to denote the bipartite graph with bipartition $(P_{i}^{f},P_{i+1}^{f})$, where, for any $x^{[k]}\in P_{i}^{f}$ and any $y^{[\ell]}\in P_{i+1}^{f}$, $x^{[k]}$ is adjacent to $y^{[\ell]}$ if and only if $y$ covers $x$. Following [19, 77.3], we call $G_{i}(P)^{f}$ the $f$-parallelization of $G_{i}(P)$ for $i=0,\ldots,r-1$.
Recall that if $G$ is a bipartite graph with partition $(V_{1},V_{2})$, a perfect matching of $G$ is defined to be a bijection $\phi$ from $V_{1}$ to $V_{2}$ such that each $a\in V_{1}$ is adjacent to its image $\phi(a)\in V_{2}$.
Lemma 3.2.
Let $f:P\rightarrow\mathbb{N}_{0}$ be a function. Then $f$ belongs to $C_{P}$ if and only if $G_{i}(P)^{f}$ has a perfect matching for all $i=0,\ldots,r-1$.
Proof.
By Definition 1.1, $f$ belongs to $C_{P}$ if and only if $U_{f}=\prod_{x\in P}x^{f(x)}\in\mathbb{K}[C_{P}]$. We define the map $\Gamma:P^{f}\rightarrow P$ such that it assigns every element of $P^{f}$ to its prototype. In other words, for any $a=x^{[j]}$, where $x\in P$ and $1\leq j\leq f(x)$, we have $\Gamma(a)=x$. It is easy to see that $U_{f}=\prod_{a\in P^{f}}\Gamma(a)$.
Suppose that for $i=0,\ldots,r-1$, the graph $G_{i}(P)^{f}$ has a perfect matching, which we denote by $\phi_{i}$. Write that
$P_{0}^{f}=\{a_{1},a_{2},\ldots,a_{n}\}.$
Let $\varphi_{i}$ denotes the composition $\phi_{i}\phi_{i-1}\cdots\phi_{0}$ for $i=0,\ldots,r-1$. Then
$P_{1}^{f}=\{\varphi_{0}(a_{1}),\varphi_{0}(a_{2}),\ldots,\varphi_{0}(a_{n})\},$
[…]
$P_{i}^{f}=\{\varphi_{i-1}(a_{1}),\varphi_{i-1}(a_{2}),\ldots,\varphi_{i-1}(a_{n})\},$
[…]
$P_{r}^{f}=\{\varphi_{r-1}(a_{1}),\varphi_{r-1}(a_{2}),\ldots,\varphi_{r-1}(a_{n})\}.$
Note that for any $1\leq j\leq n$,
$\Gamma(a_{j})<\Gamma(\varphi_{0}(a_{j}))<\cdots<\Gamma(\varphi_{i}(a_{j}))<\cdots<\Gamma(\varphi_{r-1}(a_{j}))$
is a maximal chain of $P$. Therefore, $U_{f}$ is expressed as the product
$\prod_{j=1}^{n}\Big{(}\Gamma(a_{j})\Gamma(\varphi_{0}(a_{j}))\cdots\Gamma(\varphi_{i}(a_{j}))\cdots\Gamma(\varphi_{r-1}(a_{j}))\Big{)}.$
This implies $U_{f}$ belongs to $\mathbb{K}[C_{P}]$, and so $f\in C_{P}$.
Conversely, suppose that $U_{f}$ belongs to $\mathbb{K}[C_{P}]$. Then there exist a positive integer $n\geq 1$ and $n$ maximal chains of $P$, say $x_{0,j}<x_{1,j}<\cdots<x_{r,j}$ for $j=1,\ldots,n$ such that
$U_{f}=\prod_{j=1}^{n}x_{0,j}x_{1,j}\cdots x_{r,j}.$
Denote by $K_{i}$ the multi-set $\{x_{i,1},x_{i,2},\ldots,x_{i,n}\}$ for $i=0,\ldots,r$. Note that
$U_{f}=\prod_{i=0}^{r}\prod_{x\in K_{i}}x=\prod_{i=0}^{r}\prod_{a\in P_{i}^{f}}\Gamma(a).$
Fix $i\in[0,r]$. Since $\prod_{x\in K_{i}}x=\prod_{a\in P_{i}^{f}}\Gamma(a)$, it follows that each $x\in P_{i}$ appears $f(x)$ times in $K_{i}$. Thus, there is a bijection, say $w_{i}$, from $K_{i}$ to $P_{i}^{f}$ such that $\Gamma(w_{i}(x_{i,j}))=x_{i,j}$ for all $j=1,\ldots,n$. For $i=0,\ldots,r-1$, consider the following diagram:
[…]
Here, $\mathbf{b}_{i}$ denotes the bijection $K_{i}\rightarrow K_{i+1}$ sending $x_{i,j}$ to $x_{i+1,j}$ for $j=1,\ldots,n$. Since $\mathbf{b}_{i},w_{i}$ and $w_{i+1}$ are all bijections, there is a bijection $\phi_{i}:P_{i}^{f}\rightarrow P_{i+1}^{f}$ such that the above diagram is commutative. It remains to show that $\phi_{i}$ is a perfect matching of $G_{i}(P)^{f}$. Let $a\in P_{i}^{f}$. There is a unique $1\leq j\leq n$ such that $x_{i,j}\in K_{i}$ and $w_{i}(x_{i,j})=a$. By the definition of $\phi_{i}$, $\phi_{i}(a)=w_{i+1}(x_{i+1,j}).$ Therefore, $\Gamma(a)=x_{i,j}<x_{i+1,j}=\Gamma(\phi_{i}(a))$, implying $a$ is adjacent to $\phi_{i}(a)$ in $G_{i}(P)^{f}$. Consequently, $\phi_{i}$ is a perfect matching of $G_{i}(P)^{f}$, as required.
We now move to prove Theorem 3.1.
proof of Theorem 3.1.
In this proof, the notation $G_{i}(P)^{f}$ will be written as $G_{i}^{f}$ for short.
(1$\Rightarrow$ (2) Fix $i\in[0,r-1]$. By Lemma 3.2, $G_{i}^{f}$ is a bipartite graph with partition $(P_{i}^{f},P_{i+1}^{f})$ that possesses a perfect matching. From this it follows that $|P_{i}^{f}|=|P_{i+1}^{f}|$ by Lemma 2.1. As a result, we can conclude that $f(P_{i})=f(P_{i+1})$ since $|P_{j}^{f}|=f(P_{j})$ holds for all $j=0,\ldots,r$. This establishes the validity of statement (a).
Now, let $V$ be a non-empty subset of $P_{i}$. We may assume without loss of generality that $f(V)\neq 0$, ensuring that $V^{f}$ is a non-empty subset of $P_{i}^{f}$. In view of Lemma 2.1, we deduce that $|V^{f}|\leq|N_{G_{i}^{f}}(V^{f})|$. Observing that $N_{G_{i}^{f}}(V^{f})\subseteq N_{P}(V)^{f}$ and that $f(N_{P}(V))=|N_{P}(V)^{f}|$, we arrive at the inequality $f(V)\leq f(N_{P}(V))$. Therefore, statement (b) is also proven.
(2$\Rightarrow$ (1): Fix $i\in[0,r-1]$. According to Lemma 3.2, it suffices to show that $G_{i}^{f}$ possesses a perfect matching under conditions (a) and (b). Given that $f(P_{i})=f(P_{i+1})$, we deduce that $|P_{i}^{f}|=|P_{i+1}^{f}|$. Consider any nonempty subset $S\subseteq P_{i}^{f}$. Setting $V=\Gamma(S)$, then it holds that
$N_{G_{i}^{f}}(S)=(N_{P}(V))^{f}\quad\text{and}\quad S\subseteq V^{f}.$
By virtue of (b), we have $|V^{f}|=f(V)\leq f(N_{P}(V))$, whence it follows that
$|S|\leq|V^{f}|=f(V)\leq|(N_{P}(V))^{f}|=|N_{G_{i}^{f}}(S)|.$
Consequently, Lemma 2.1 implies that $G_{i}^{f}$ indeed has a perfect matching.
Now, we are ready to prove that the chain algebra $\mathbb{K}[C_{P}]$ is normal.
Corollary 3.3.
Let $P$ be a finite pure poset. Then $C_{P}$ is a normal affine semigroup. Consequently, $\mathbb{K}[C_{P}]$ is a normal domain.
Proof.
Let $f,g,$ and $h$ be functions belonging to $C_{P}$ such that $nf=ng+h$ for some $n>0$. To prove that $C_{P}$ is normal, it suffices to show that $h=nh^{\prime}$ for some $h^{\prime}\in C_{P}$. Define $\alpha:=g-f$. For any $x\in P$, we have $\alpha(x)=g(x)-f(x)$ is an integer. Given that $n\alpha(x)=h(x)$ and $h(x)$ is non-negative, it follows that $\alpha(x)\geq 0$. This establishes that $\alpha$ is a function from $P$ to $\mathbb{N}_{0}$. Furthermore, since $\alpha=\frac{h}{n}$ and $h\in C_{P}$, it holds that for all pure subsets $X$ of $P$, $\alpha(X)\leq\alpha(N_{P}(X))$. Additionally, for all $0\leq i,j\leq r$, we have $\alpha(P_{i})=\alpha(P_{j})$. These properties of $\alpha$ imply $\alpha\in C_{P}$ by Theorem 3.1.
Therefore, we have $h=n\alpha$ and $\alpha\in C_{P}$, and thus $C_{P}$ is a normal semigroup. Finally, the normality of $\mathbb{K}[C_{P}]$ follows from the normality of $C_{P}$.
4. Chain polytope of $P$
In this section, we prove the chain polytope of any finite pure poset has the integer decomposition property, and characterize when it is indecomposable.
According to Theorem 3.1, $C_{P}$ is a homogenous affine semigroup wherein $\deg(f)=f(P_{i})$ for any $i\in[0,r]$. This implies that any $f\in(C_{P})_{k}$ is the sum of $k$ functions from $(C_{P})_{1}$, where $k$ is any integer $>0$. It is worth noting that $(C_{P})_{1}$ is the set of minimal generators of $C_{P}$. Here, $(C_{P})_{k}$ is the set of the functions in $C_{P}$ of degree $k$.
Proposition 4.1.
Let $P$ be a finite pure poset. Then $D_{P}$ has the integer decomposition property. Consequently, $\mathbb{K}[C_{P}]$ is the same as the Ehrhart ring $A(D_{P})$.
Proof.
Let $f\in kD_{P}\cap\mathbb{Z}^{P}$, where $k$ is any positive integer. To prove that $D_{P}$ has IDP, it suffices to show that $f$ is contained in $(C_{P})_{k}$. We do this in the following steps. First, we notice that that $f$ is a function from $P$ to $\mathbb{N}_{0}$. Next, by Definition 1.2, we may write $f=k(a_{1}f_{1}+\cdots+a_{t}f_{t})$, where $f_{i}\in(C_{P})_{1}$ for all $i=1,\ldots,t$ and the coefficients $a_{i}$ satisfy $a_{1}+\cdots+a_{t}=1$ with $a_{i}\geq 0$ for all $i=1,\ldots,t$. According to Theorem 3.1, for all $i=1,\ldots,t$, we have $f_{i}(P_{j})=1$ for $j=0,\ldots,r$ and $f_{i}(X)\leq f_{i}(N_{P}(X))$ for all pure set $X\subseteq P$. From these, it follows that $f(P_{j})=k$ for $j=0,\ldots,r$ and that
$f(X)=k\sum\limits_{j=1}^{t}a_{j}f_{j}(X)\leq k\sum\limits_{j=1}^{t}a_{j}f_{j}(N_{P}(X))=f(N_{P}(X))$
for all pure set $X\subseteq P$. Thus, $f$ belongs to $(C_{P})_{k}$, as required.
By [9, Theorem 4.3], if a lattice polytope $D$ has the integer decomposition property, then $\mathbb{K}[D]$ is normal. However, the converse is not true in general, as shown by [9, Example 4.4]. Hence, Corollary 3.3 can also be derived from Proposition 4.1 together with [9, Theorem 4.3]. We now classify when $D_{P}$ is indecomposable.
Theorem 4.2.
Let $P$ be a finite pure poset with rank $r\geq 1$. Then the following statements are equivalent:
  1. (1)
    The chain polytope $D_{P}$ is indecomposable.
  2. (2)
    $G_{i}(P)$ is not a complete bipartite graph for all $i=0,\ldots,r-1$.
  3. (3)
    $P$ cannot be expressed as the ordinal sum of two non-empty sub-posets of $P$.
Proof.
$(2)\Rightarrow(3)$ Suppose that $P$ is the ordinal sum of non-empty sub-posets $Q_{1}$ and $Q_{2}$ of $P$. Since $P$ is pure, it follows that both $Q_{1}$ and $Q_{2}$ are pure. Let $k:=\max\{\rho(a)\colon\;a\in Q_{1}\}$. Then $G_{k}(P)$ is a complete bipartite graph, a contradiction.
$(3)\Rightarrow(2)$ If $G_{i}(P)$ is a complete bipartite graph for some $i$, then $P$ is the ordinal sum of $Q_{1}:=P_{0}\cup\cdots\cup P_{i}$ and $Q_{2}:=P_{i+1}\cup\cdots\cup P_{r}$, a contradiction.
If $P$ is the ordinal sum $Q_{1}\oplus Q_{2}$, then $D_{P}=D_{Q_{1}}\times D_{Q_{2}}$. This proves $(3)\Rightarrow(1)$. It remains to prove $(2)\Rightarrow(1)$.
$(2)\Rightarrow(1)$ Suppose on the contrary that $D_{P}$ is not indecomposable. Then, according to [6, Lemma 30], there exist $(0,1)$-polytopes $D_{1}\subset\mathbb{R}^{Q_{1}}$ and $D_{2}\subset\mathbb{R}^{Q_{2}}$ such that $D_{P}$ is the direct product $D_{1}\times D_{2}$. Here, $Q_{1}$ and $Q_{2}$ are non-empty subsets of $P$ such that $Q_{1}\cap Q_{2}=\emptyset$ and $Q_{1}\cup Q_{2}=P$.
We use the support of $f$ to refer to the subset $\{x\in P\colon\;f(x)=1\}$ for any $f\in D_{P}\cap\mathbb{Z}^{P}$ or $f\in D_{1}\cap\mathbb{Z}^{Q_{1}}$ or $f\in D_{2}\cap\mathbb{Z}^{Q_{2}}$. It is clear that every $f$ is determined by its support. By $D_{P}=D_{1}\times D_{2}$, it means that if $S$ is the support of a $(0,1)$-vector in $D_{P}$, then $S\cap Q_{1}$ and $S\cap Q_{2}$ are the supports of $(0,1)$-vectors in $D_{1}$ and $D_{2}$ respectively. Conversely, if $S_{1}$ and $S_{2}$ are the supports of $(0,1)$-vectors in $D_{1}$ and $D_{2}$ respectively, then $S_{1}\cup S_{2}$ is the support of a $(0,1)$-vector in $D_{P}$, i.e., a maximal chain of $P$.
We claim that for any $a_{1}\in Q_{1}$ and any $a_{2}\in Q_{2}$, $a_{1}$ and $a_{2}$ are comparable in $P$. Since $P$ is pure, there exist maximal chains $C_{1}$ and $C_{2}$ of $P$ such that $a_{j}\in C_{j}$ for $j=1,2$. Since $Q_{j}\cap C_{1}$ and $Q_{j}\cap C_{2}$ are supports of some (0,1)-vectors of $D_{j}$ for $j=1,2$, the union of $Q_{1}\cap C_{1}$ and $Q_{2}\cap C_{2}$ is a maximal chain of $P$. Given that $a_{i}\in Q_{i}\cap C_{i}$ for $i=1,2$, it follows that $a_{1}$ and $a_{2}$ belongs to a maximal chain of $P$. In particular, $a_{1}$ and $a_{2}$ are comparable, proving the claim.
By this claim, we have for each $i=0,\ldots,r$, either $P_{i}\subseteq Q_{1}$ or $P_{i}\subseteq Q_{2}$. Since both $Q_{1}$ and $Q_{2}$ are non-empty, there exists $k\in[0,r-1]$ such that $P_{k}\subseteq Q_{1}$ and $P_{k+1}\subseteq Q_{2}$. However, the claim implies that $G_{k}(P)$ is a complete bipartite graph, a contradiction.
According to [6, Theorem 32], if $D$ is an indecomposable lattice polytope, then the Enrhart ring $A(D)$ is Gorenstein if and only if it is nearly Gorenstein. The following example illustrates an application of Theorem 4.2.
Example 4.3.
Let $G$ be a connected bipartite graph. Note that a bipartite graph can be regarded as a pure poset of rank one, and its edge algebra is isomorphic to the chain algebra of the corresponding poset. By Theorem 4.2, the edge polytope of $G$ is indecomposable precisely when $G$ is not complete. Hence, according to [6, Theorem 32], the edge ring $\mathbb{K}[G]$ of $G$ is Gorenstein if and only if it is nearly Gorenstein in the case when $G$ is not complete.
5. $K_{P}$ and Canonical modules
In this section we provide an explicit description of the canonical module of $\mathbb{K}[C_{P}]$ for every pure poset $P$. To do so, the following definition is essential.
Definition 5.1.
Let $P$ be a pure poset. A subset of $P$ is pure if it consists of elements of same rank. For a pure subset $X$ of $P$, we call $X$ to be strict if there is $f\in C_{P}$ such that $f(X)<f(N_{P}(X))$.
Recall that $K_{P}$ is defined as the ideal of $C_{P}$ consisting of all functions $f\in C_{P}$ such that for any $g\in C_{P}$, there exist a positive integer $n$ and an element $h$ in $C_{P}$ satisfying the equation $nf=g+h$. By Lemma 5.5, the canonical module of $\mathbb{K}[C_{P}]$ is isomorphic to the ideal of $\mathbb{K}[C_{P}]$ generated by $\prod_{x\in P}x^{f(x)}$ with $f\in K_{P}$. Hence, elucidating the structure of the canonical module of $\mathbb{K}[C_{P}]$ is equivalent to characterizing the functions contained in $K_{P}$.
Lemma 5.2.
Let $f\in C_{P}$. Then $f$ belongs to $K_{P}$ if and only if the following conditions hold:
$\mathrm{(1)}$ $f(X)<f(N_{P}(X))$ for any strict subset $X$ of $P$;
$\mathrm{(2)}$ $f(x)>0$ for all $x\in P$.
Proof.
$\Rightarrow$ Let $f$ be a function belonging to $K_{P}$. Consider any strict subset $X$ of $P$. There exists $g\in C_{P}$ such that $g(X)<g(N_{P}(X))$. Given that $f\in K_{P}$, it follows that there exists $n>0$ and $h\in C_{P}$ such that $nf=g+h$. As a result, we have
$f(N_{P}(X))-f(X)=\frac{1}{n}(g(N_{P}(X))-g(X))+\frac{1}{n}(h(N_{P}(X))-h(X))>0.$
Take an arbitrary element $x\in P$. Since there is at least one maximal chain of $P$ containing $x$, it follows that there is $g^{\prime}\in C_{P}$ such that $g^{\prime}(x)>0$. Since $nf=g^{\prime}+h^{\prime}$ for some $n>0$ and for some $h^{\prime}\in C_{P}$, we obtain $nf(x)=g^{\prime}(x)+h^{\prime}(x)>0$. From this, we conclude that $f(x)>0$.
$\Leftarrow$ Suppose that $f$ is a function in $C_{P}$ that fulfills both conditions (1) and (2). Given an arbitrary function $g\in C_{P}$, it is routine to verify that the function $h:=nf-g$ belongs to $C_{P}$, where $n$ is given as
$n:=\max\{g(N_{P}(X))-g(X)\colon\;X\mbox{ is a strict set of }P\}+\max\{g(x)\colon\;x\in P\}.$
Therefore, we conclude that $f\in K_{P}$, as desired.
We next identify all the strict subsets of $P$. To this end, we introduce several additional notions.
Fix $i\in\{0,\ldots,r-1\}$. Then the bipartite graph $G_{i}(P)$ is the disjoint union of its connected components, say $G_{i,1},\ldots,G_{i,k_{i}}$. Note that all $G_{i,j}$ are connected bipartite graphs and we may assume that $G_{i,j}$ has a bipartition $(P_{i,j},N_{P}(P_{i,j}))$ such that $P_{i,j}\subseteq P_{i}$ for $j=1,\ldots,k_{i}$. One easily observes that $P_{i}=P_{i,1}\sqcup\cdots\sqcup P_{i,k_{i}}$ and $P_{i+1}=N_{P}(P_{i,1})\sqcup\cdots\sqcup N_{P}(P_{i,k_{i}}).$
Definition 5.3.
For all $i=0,\ldots,r-1$ and $j=1,\ldots,k_{i}$, we define $P_{i,j}$ as a connected component of rank $i$.
It is evident that every connected component of rank $i$ is a pure subset of $P$.
Lemma 5.4.
Let $P$ be a pure poset of rank $r$, and fix an integer $i$ with $10\leq i\leq r-$ 1. Then the following statements hold:
$\mathrm{(1)}$ Every connected component of rank $i$ is not a strict set.
$\mathrm{(2)}$ Every non-empty proper subset of any connected component of rank $i$ is a strict set.
$\mathrm{(3)}$ Consider any pure subset $X\subseteq P_{i}$. The set $X$ is a strict set if and only if there exists at least one $j\in\{1,\ldots,k_{i}\}$ such that $X\cap P_{i,j}$ is a non-empty proper subset of $P_{i,j}$.
Proof.
(1) Fix $f\in C_{P}$. Then $f(P_{i,j})\leq f(N_{P}(P_{i,j}))$ for all $j=1,\ldots,k_{i}$. Furthermore, $f(P_{i})=\sum_{j=1}^{k_{i}}f(P_{i,j})$ and $f(P_{i+1})=\sum_{j=1}^{k_{i}}f(N_{P}({P_{i,j}}))$. Note that $f(P_{i})=f(P_{i+1})$, it follows that $f(P_{i,j})=f(N_{P}(P_{i,j}))$ for all $j=1,\ldots,k_{i}$. Thus, every $P_{i,j}$ is not a strict set.
(2) Let $S$ be a non-empty proper subset of $P_{i,j}$ for some $i,j$. Since the bipartite graph $G_{i,j}$ with bipartition $(P_{i,j},N_{P}(P_{i,j}))$ is connected, there exist $\alpha\in S$ and $\beta\in P_{i,j}\setminus S$ such that $N_{P}(\alpha)\cap N_{P}(\beta)$ is non-empty. Let $\gamma\in N_{P}(\alpha)\cap N_{P}(\beta)$. Then there is at least one maximal chain of the following form:
$\alpha_{0}<\alpha_{1}<\alpha_{2}<\cdots<\alpha_{i-1}<\beta<\gamma<\gamma_{i+1}<\cdots<\gamma_{r}.$
Denote by $w$ the function from $P$ to $\mathbb{N}_{0}$ which corresponds to this chain. That is $w(p)=1$ if $p$ belongs to this chain and $w(p)=0$ otherwise. Since $\gamma$ belongs to $N_{P}(S)$, it follows that
$\sum_{p\in S}w(p)=0\mbox{\qquad and \qquad}\sum_{p\in N_{P}(S)}w(p)\geq w(\gamma)=1.$
It particularly implies $w(S)<w(N_{P}(S))$, thereby proving $S$ is a strict set.
(3) Set $X_{j}=X\cap P_{i,j}$. Then $X=\sqcup_{j=1}^{k_{i}}X_{j}$ and $N_{P}(X)=\sqcup_{j=1}^{k_{i}}N_{P}(X_{j})$. This implies $f(X)<f(N_{P}(X))$ if and only if there is at least one $j\in[1,k_{i}]$ such that $f(X_{j})<f(N_{P}(X_{j}))$. From this, the result follows.
We are now in the position to present the main result of this section.
Theorem 5.5.
Let $P$ be a pure poset of rank $r$. Then for any $f\in C_{P}$, $f$ belongs to $K_{P}$ if and only if the following conditions are satisfied:
$\mathrm{(1)}$ For any $10\leq i\leq r-$ 1 and for any non-empty proper subset of any connected component of rank $i$, it holds that $f(X)<f(N_{P}(X))$.
$\mathrm{(2)}$ For all $x\in P$, $f(x)>0$
Proof.
Let $f\in C_{P}$. Given Lemmas 5.2 as well as 5.4, we only need to show that if condition (1) holds, then $f(X)<f(N_{P}(X))$ for any strict subset $X$ of $P$. Let $X\subseteq P_{i}$ be a strict subset of $P$. By Lemma 5.4.(3), there exists $1\leq j_{0}\leq k_{i}$ such that $X\cap P_{i,j_{0}}$ is a non-empty proper subset of $P_{i,j_{0}}$. It follows from condition (1) that $f(X\cap P_{i,j_{0}})<f(N_{P}(X\cap P_{i,j_{0}}))$. Consequently, we have
$f(X)=\sum\limits_{j=1}^{k_{i}}f(X\cap P_{i,j})<\sum\limits_{j=1}^{k_{i}}f(N_{P}(X\cap P_{i,j}))=f(N_{P}(X)),$
as required.
6. Dimension of a chain algebra
In this section, we present an explicit formula for the Krull dimension of chain algebras, alongside a definitive formula for the dimension of chain polytopes associated with pure posets. We begin with a result that may be well-known in linear algebra; however, for completeness, we include a proof.
Lemma 6.1.
Let $V$ be a linear subspace of the $\mathbb{Q}$-space $\mathbb{Q}^{n}$ and let $S$ be an open subset of $\mathbb{Q}^{n}$. If $S\cap V\neq\emptyset$, then $V$ is spanned by the vectors belonging to $S\cap V$.
Proof.
Let $\alpha_{0}\in S\cap V$ and let $\alpha_{1},\ldots,\alpha_{r}$ be a $\mathbb{Q}$-basis for $V$. Since $\alpha_{0}$ is an interior point of $S$, there exists a positive integer $k>0$ such that $\frac{1}{k}(\alpha_{i}-\alpha_{0})+\alpha_{0}$ belongs to $S$ for $i=1,\ldots,r$. Thus, $S\cap V$ contains the vectors $\frac{1}{k}(\alpha_{i}-\alpha_{0})+\alpha_{0}$ for $i=1,\ldots,r$ as well as $\alpha_{0}$. This implies that $\alpha_{1},\ldots,\alpha_{r}$ belong to $\langle S\cap V\rangle_{\mathbb{Q}}$. Here, $\langle S\cap V\rangle_{\mathbb{Q}}$ denotes the subspace of $\mathbb{Q}^{n}$ spanned by $S\cap V$. Hence, we have $V=\langle S\cap V\rangle_{\mathbb{Q}}$, as required.
We refer to Definition 5.3 for the definition of connected components of rank $i$.
Theorem 6.2.
Let $P$ be a finite pure poset with rank r. Let $k_{i}$ denote the number of connected components of rank $i$ for $i=0,\ldots,r-1$. Then
$\dim\mathbb{K}[C_{P}]=|P|-(k_{0}+\cdots+k_{r-1}).$
Proof.
For each $p\in P$, let $X_{p}$ denote an indeterminate indexed by $p$. Consider the following systems of equalities or inequalities:
(§) $\sum_{p\in P_{i,j}}X_{p}=\sum_{p\in N_{P}(P_{i,j})}X_{p},\qquad\mbox{ for }i=0,\ldots,r-1,j=1,\ldots,k_{i}.$
(†) $X_{p}\geq 0,\quad\mbox{ for }p\in P.$
(${\dagger}_{<}$) $X_{p}>0,\quad\mbox{ for }p\in P.$
(‡) $\sum_{p\in S}X_{p}\leq\sum_{p\in N_{P}(S)}X_{p},\mbox{ for }\emptyset\neq S\subsetneq P_{i,j},1\leq i\leq r-1,1\leq j\leq k_{i}.$
(${\ddagger}_{<}$) $\sum_{p\in S}X_{p}<\sum_{p\in N_{P}(S)}X_{p},\mbox{ for }\emptyset\neq S\subsetneq P_{i,j},1\leq i\leq r-1,1\leq j\leq k_{i}.$
It is natural to identify every rational or integral solution to part or all of these equalities or inequalities with a function mapping from $P$ to the set of rational numbers or integral numbers. Under this identification, we have
[…]
  • $K_{P}$ consists of all integral solutions that satisfy $(\S),({\dagger}_{<})$ and $({\ddagger}_{<})$.
For a subset $M$ of the $\mathbb{Q}$-subspace $\mathbb{Q}^{P}$ (the set of all functions from $P$ to $\mathbb{Q}$), we denote by $\langle M\rangle_{\mathbb{Q}}$ the $\mathbb{Q}$-subspace of $\mathbb{Q}^{P}$ spanned by $M$. It is evident that the Krull dimension of $\mathbb{K}[C_{P}]$ is equal to $\dim_{\mathbb{Q}}\langle C_{P}\rangle_{\mathbb{Q}}$. We also introduce the following notions:
  • Let $V$ denote the set of all rational solutions to (§)§($\S$)( § );
  • Let $A$ denote the set of all rational solutions to $({\dagger}_{<})$ and $({\ddagger}_{<})$;
[…]
Note that $V$ is a $\mathbb{Q}$-linear subspace of $\mathbb{Q}^{P}$, while $A$ is an open subset of $\mathbb{Q}^{P}$. It is straightforward to observe that $V\cap A\subseteq V\cap B\subseteq V$. Since $K_{P}\subseteq V\cap A$, it follows that $V\cap A$ is non-empty. Therefore, according to Lemma 6.1, $\langle V\cap A\rangle_{\mathbb{Q}}$ coincides with $V$. Given that $\langle C_{P}\rangle_{\mathbb{Q}}=\langle V\cap B\rangle_{\mathbb{Q}}$, it follows that the Krull dimension of $\mathbb{K}[C_{P}]$ is equal to the dimension $\dim_{\mathbb{Q}}V$. Finally, observing that the rows of the coefficient matrix of the homogeneous system (§)§($\S$)( § ) of linear equations are linear independent (in fact, the coefficient matrix of (§)§($\S$)( § ) is in row echelon form), the result follows from the basic linear algebra.
An application of Theorem 6.2 recovers [4, Theorem 3.4].
Corollary 6.3.
Let $L$ be a finite distributive lattice of rank $r$. Then the chain algebra $\mathbb{K}[C_{L}]$ has the dimension $|L|-r$.
Proof.
In view of the initial portion of the proof of [4, Theorem 3.4], $G_{i}(L)$ is a connected bipartite graph for each $10\leq i\leq r-$ 1. Consequently, the desired result follows from Theorem 6.2.
In view of [9, Section 4.2.4] together with [9, Theorem 3.1], the dimension of a lattice polytope $D$ is equal to the dimension of the toric ring $\mathbb{K}[D]$ minus one. Hence, the following result is immediate.
Corollary 6.4.
With the same assumptions as in Theorem 6.2, we have $\dim D_{P}=|P|-(k_{0}+\cdots+k_{r-1})-1$.
7. Posets of width 2
Recall the width of a poset is the maximal size of an anti-chain contained in this poset. Let $P$ be a pure poset of width 2. In this section, we will compute the regularity of $\mathbb{K}[C_{P}]$ and characterize the conditions under which $\mathbb{K}[C_{P}]$ is Gorenstein or nearly Gorenstein.
Hereafter, we always assume that $P$ is a pure poset of width 2 and rank $r\geq 1$ unless otherwise stated. As previously, we denote by $P_{i}$ the set of elements of $P$ with rank $i$, and by $G_{i}(P)$ the bipartite graph with bipartition $(P_{i},P_{i+1})$, where for $a\in P_{i}$ and $b\in P_{i+1}$, $a$ is adjacent to $b$ if and only if $a<b$. It is clear that $P_{i}$ contains at most 2 elements for each $i$. We make the following observations:
(1) If, for some $i$, $G_{i}(P)$ is not connected, then $G_{i}(P)$ consists of two disjoint edges. We may assume $P_{i}=\{a_{1},a_{2}\}$ and $P_{i+1}=\{b_{1},b_{2}\}$ with $a_{j}<b_{j}$ for $j=1,2$. In this case, we have $f(a_{i})=f(b_{i})$ for $i=1,2$ and for $f\in C_{P}$. Thus, if we identify $a_{j}$ with the $b_{j}$ for $j=1,2$ to obtain a smaller poset $P^{\prime}$, then it is evident that $C_{P}$ is isomorphic to $C_{P^{\prime}}$, and $K_{P}$ is isomorphic to $K_{P^{\prime}}$. In particular, we have $\mathbb{K}[C_{P}]$ is isomorphic to $\mathbb{K}[C_{P^{\prime}}]$. Hereafter, we denote by $\overline{P}$ the poset that is obtained from $P$ by a series of such identifications, where $G_{i}(\overline{P})$ are connected for all possible $i$. For example, if $P$ is the disjoint union of two chains of the same length, then $\overline{P}$ is an anti-chain containing exactly two elements.
(2) If there is an index $0<i<r$ such that $G_{i}(P)$ is complete bipartite, then $P$ is the ordinal sum of the subposets $Q_{1}:=P_{0}\cup\cdots\cup P_{i}$ and $Q_{2}:=P_{i+1}\cup\cdots\cup P_{r}$. Furthermore, if $|P_{i}|=1$, then $\mathbb{K}[C_{Q_{1}}]$ is isomorphic to $\mathbb{K}[C_{Q_{1}^{\prime}}]$, where $Q_{1}^{\prime}$ denotes the subposet $Q_{1}\setminus P_{i}$, and similarly, if $|P_{i+1}|=1$, then $\mathbb{K}[C_{Q_{2}}]$ is isomorphic to $\mathbb{K}[C_{Q_{2}^{\prime}}]$, where $Q_{2}^{\prime}:=Q_{2}\setminus P_{i+1}$.
The following definition is now natural.
Definition 7.1.
A pure poset $P$ is called basic if it is of wideth 2 and rank $\geq 1$, and $G_{i}(P)$ is a connected bipartite graph that is not complete for each $i=0,\ldots,r-1$.
The above discussions establish that, given any pure poset $P$ with width 2, we have a poset $\overline{P}$ such that $G_{i}(\overline{P})$ is connected for all possible $i$ and $\mathbb{K}[C_{P}]$ is isomorphic to $\mathbb{K}[C_{\overline{P}}]$. Furthermore, $\overline{P}$ is the ordinal sum of basic posets and posets of rank 00 and width 2. Note that a poset of rank 00 and width 2 is nothing but an anti-chain containing two elements.
In what follows, whenever $P$ is a basic poset, we always assume that $P=\{X_{0},\ldots,X_{r},Y_{0},\ldots,Y_{r}\}$ such that $X_{0}<X_{1}<\cdots<X_{r}$ and $Y_{0}<\cdots<Y_{r}$ are two maximal chains of $P$. For $X\in P$, we say $X$ has degree $i$ if $|N_{P}(X)|=i$, meaning that there are exactly $i$ elements that covers $X$. By our assumption, for each $i\in[0,r-1]$, exactly one of $\{X_{i},Y_{i}\}$ has degree one. Furthermore, if $X_{i}$ has degree one, then $N_{P}(X_{i})=\{X_{i+1}\}$ and $N_{P}(Y_{i})=\{X_{i+1},Y_{i+1}\}$, and similarly, if $Y_{i}$ has degree one then $N_{P}(Y_{i})=\{Y_{i+1}\}$ and $N_{P}(X_{i})=\{X_{i+1},Y_{i+1}\}$. Thus, To fully describe a basic poset, it is enough to identify all the elements that have degree one. We associate every basic poset a sequence of positive integers as follows.
Definition 7.2.
Let $(c_{1},c_{2},\ldots,c_{n})$ be an integral vector, where $c_{1}+\cdots+c_{n}=r$ with $c_{i}\geq 1$ for all $i=1,\ldots,n$. Set $b_{0}=0$ and $b_{i}=c_{1}+c_{2}+\cdots+c_{i}$ for $i=1,\ldots,n$. Let $P$ be a basic poset. By relabelling if necessary, we may assume $X_{0}$ has degree one. We say a basic poset $P$ belongs to type $(c_{1},c_{2},\ldots,c_{n})$ if the degree-one elements of $P$ are as follows:
$X_{0},\ldots,X_{c_{1}-1},Y_{c_{1}},\ldots,Y_{c_{1}+c_{2}-1},\ldots,$
$X_{b_{2i}},X_{b_{2i}+1},\ldots,X_{b_{2i+1}-1},Y_{b_{2i+1}},Y_{b_{2i+1}+1},\ldots,Y_{b_{2i+2}-1},\ldots$
$W_{b_{n-1}},W_{b_{n-1}+1},\ldots,W_{b_{n}-1}.$
Here, $W=X$ if $n$ is odd, and $W=Y$ if $n$ is even.
Up to isomorphisms, every basic poset is determined by its type. Specializations of Theorems 3.1 and 5.5 to basic posets produces the following result.
Lemma 7.3.
Let $P$ be a basic poset of type $(c_{1},\ldots,c_{n})$ with $c_{0}$ set to 0. Consider a function $f:P\rightarrow\mathbb{N}_{0}$ such that $f(P_{i})=f(P_{j})$ for all $0\leq i,j\leq r$. Then
[…]
  • If $i$ is even, then $f(X_{k})\leq f(X_{k+1})$;
  • If $i$ is odd, then $f(X_{k})\geq f(X_{k+1})$.
[…]
$f(X_{0})\leq\cdots\leq f(X_{c_{1}})\geq f(X_{c_{1}+1})\geq\cdots\geq f(X_{c_{1}+c_{2}})\leq\cdots\geq\cdots f(X_{r-1})?f(X_{r}).$
Here, ???? is $\leq$ if $n$ is odd and ???? is $\geq$ if $n$ is even.
[…]
  • If $i$ is even, then $f(X_{k})<f(X_{k+1})$;
  • If $i$ is odd, then $f(X_{k})>f(X_{k+1})$.
[…]
$f(X_{0})<\cdots<f(X_{c_{1}})>f(X_{c_{1}+1})>\cdots>f(X_{c_{1}+c_{2}})<\cdots>\ldots f(X_{r-1})?f(X_{r}).$
Here, ???? is $<$ if $n$ is odd and ???? is $>$ if $n$ is even.
Henceforth, we shall use $\lambda(P)$ (or simply $\lambda$) to represent the maximum value among $c_{i}$ for $1\leq i\leq n$. Additionally, we set $b_{0}=0$, and $b_{i}=c_{1}+\ldots+c_{i}$ for $i=1,\ldots,r$. Recall that for a function $f\in C_{P}$, the degree of $f$, denoted by $\deg(f)$, is defined as $f(P_{i})$, where $i$ can be any number within $[0,r]$. The following lemma allows us to construct functions in $K_{P}$ satisfying certain properties conveniently.
Lemma 7.4.
[…]
Proof.
We extend the definition of $f$ in the following way: For $0\leq i<s$ or $t\leq i\leq n$,
  • If $i$ is even, we define $f(X_{b_{i}})=1,f(X_{b_{i}+1})=2,\ldots,f(X_{b_{i+1}-1})=c_{i}$.
  • If $i$ is odd, we define $f(X_{b_{i}})=\lambda+1$ when $i\neq t$. It is worth noting that if $i=t$, we do not need to define $f(X_{b_{i}})$. Furthermore, we specify the values of $f(X_{b_{i}+1}),\ldots,f(X_{b_{i+1}-1})$ in such a way that they form a strictly decreasing sequence within the interval $[2,\lambda]$. As an example, such a sequence could be $\lambda,\lambda-1,\ldots,\lambda-c_{i}+2$.
  • For all $0\leq i\leq r$, we define $f(Y_{i})=d-f(X_{i})$.
Note that if either $c_{i+1}=1$ or $i=n$, we do not need to define $f(X_{b_{i}+1}),\ldots,f(X_{b_{i+1}-1})$.
By applying Lemma 7.3, it can be easily verified that $f$ indeed belongs to $K_{P}$ and has a degree of $d$, thereby completing the proof.
We now evaluate the a-invariant of the chain algebra of a basic poset.
Lemma 7.5.
Let $P$ be a basic poset of type $(c_{1},\ldots,c_{n})$. Then $a(\mathbb{K}[C_{P}])=-\lambda-2$.
Proof.
We first show that $\deg(f)\geq\lambda+2$ for all $f\in K_{P}$. Let $f\in K_{P}$. There exists $1\leq i\leq n$ such that $c_{i}=\lambda$. We may assume that $i$ is even w.l.o.g. Set $b_{0}=0$ and $b_{j}=c_{1}+\cdots+c_{j}$ for $j=1,\ldots,n$. Then
$1\leq f(X_{b_{i}})<\cdots<f(X_{b_{i}+c_{i}}).$
This implies $f(X_{b_{i}+c_{i}})\geq\lambda+1$ and so $\deg(f)=f(X_{b_{i}+c_{i}})+f(Y_{b_{i}+c_{i}})\geq\lambda+2$.
By Lemma 7.4, there is a function in $K_{P}$, say $g$, such that $g(X_{b_{i}+j})=1+j$ for $j=0,\ldots,c_{i}=\lambda$ with degree $\lambda+2$. This completes the proof.
Based on this result, we can compute the regularity of the chain algebra of all pure posets of width 2. Recall that the Hilbert series of the chain algebra $\mathbb{K}[C_{P}]$ can be expressed as a rational function $\frac{Q(t)}{(1-t)^{d}}$, where $d$ is equal to the Krull dimension of $\mathbb{K}[C_{P}]$. Since $\mathbb{K}[C_{P}]$ is Cohen-Macaulay, the regularity of $\mathbb{K}[C_{P}]$ is equal to the degree of $Q(t)$. Consequently, we have $\mathrm{reg}(\mathbb{K}[C_{P}])=d+a(\mathbb{K}[C_{P}])$. The dimension $d$ can be determined using Theorem 6.2. As stated in Subsection 2.6, the $a$-invariant of the chain algebra of the ordinal sum of two posets $P_{1}$ and $P_{2}$ is given by $a(\mathbb{K}[C_{P_{1}\oplus P_{2}}])=\min\{a(\mathbb{K}[C_{P_{1}}]),a(\mathbb{K}[C_{P_{2}}])\}.$ It is known that every pure poset of width 2 can be expressed as the ordinal sum of basic posets and posets of rank 0 and width 2. If $P$ is a poset of rank 0 and width 2, it is straightforward to see that $\mathbb{K}[C_{P}]$ is Gorenstein with an $a$-invariant of $-2$. Thus, for pure posets of width 2, the a-invariant and hence the regularity of $\mathbb{K}[C_{P}]$ can be computed using these observations and Lemma 7.5.
Finally, we will characterize when $\mathbb{K}[C_{P}]$ is Gorenstein or nearly Gorenstein whenever $P$ is a pure poset of width 2. Note that for any $f\in C_{P}$, $\deg(f)$ is equal to $f(P_{i})$ for any $i\in[0,r]$.
Lemma 7.6.
Let $P$ be a simple of type $(c_{1},\ldots,c_{n})$. Then the following statements are equivalent:
  1. (1)
    $\mathbb{K}[C_{P}]$ is pseudo-Gorenstein.
  2. (2)
    There exists a unique function belonging to $K_{P}$ with degree $\lambda+2$.
  3. (3)
    It holds that $c_{i}\in\{1,\lambda\}$ for all $i=1,\ldots,n$, and that if $c_{i}=1$ then $1<i<n$ and $c_{i-1}=c_{i+1}=\lambda$.
Proof.
$(1)\Leftrightarrow(2)$ By Lemma 7.5, we have $a(\mathbb{K}[C_{P}])=-\lambda-2$. Due to this fact, the equivalence of (1) and (2) follows by the definition of pseudo-Gorensteinness.
$(3)\Rightarrow(2)$ Let $f$ be an arbitrary function in $K_{P}$ with $\deg(f)=\lambda+2$. To establish the result, it suffices to prove that the values of $f(X)$ for all $X\in P$ are uniquely determined by the sequence $c_{1},\ldots,c_{n}$. We proceed through the following steps:
Step 1: Given that $c_{1}=\lambda$, we invoke Lemma 7.3 to conclude:
$11\leq f(X_{0})<f(X_{1})<\cdots<f(X_{\lambda})\leq\lambda+$ 1 .
Consequently, it must be that $f(X_{j})=j+1$ for $j=0,\ldots,\lambda$.
Step 2: Consider now the case where $c_{i+1}=\lambda$ for some $i$. Utilizing a similar argument as in Step 1, we deduce:
  • If $i$ is even, then $f(X_{b_{i}+j})=j+1$ for $j=0,\ldots,\lambda$.
  • If $i$ is odd, then $f(X_{b_{i}+j})=\lambda+1-j$ for $j=0,\ldots,\lambda$.
In particular, for each $i$ such that $c_{i+1}=\lambda$, the values of $f$ on the subset
$\{X_{b_{i}},X_{b_{i}+1},\ldots,X_{b_{i}+c_{i+1}-1},X_{b_{i+1}}\}$
of $P$ are uniquely specified.
Step 3: For indices $i$ where $c_{i+1}=1$, note that $c_{i}=c_{i+2}=\lambda$. By the results of Step 2, the values of $f$ at $X_{b_{i}}$ and $X_{b_{i+1}}$ are already uniquely determined. Therefore, the values of $f(X_{i})$ for $i=0,\ldots,r$ are uniquely determined. Since $f(X_{i})+f(Y_{i})=\lambda+2$, it follows that the value of $f(Y_{i})$ is also uniquely determined. In conclusion, the values of $f(X)$ for all $X\in P$ are uniquely specified, completing the proof.
$(2)\Rightarrow(3)$ Suppose on the contrary that there is $1\leq i\leq n$ such that $1<c_{i}<\lambda$. If $i$ is odd, we define functions $f$ and $g$ on the subset $\{X_{b_{i-1}},X_{b_{i-1}+1}\ldots,X_{b_{i-1}+c_{i}-1},X_{b_{i}}\}$ of $P$ by $f(X_{b_{i+1}+j})=j+1$ for $j=0,1,\ldots,c_{i}-1$ and $f(b_{i})=\lambda+1$, and $g(X_{b_{i+1}})=1$, $g(X_{b_{i+1}+j})=j+2$ for $j=1,\ldots,c_{i}-1$ and $g(b_{i})=\lambda+1$. By Lemma 7.4, both $f$ and $g$ could be extended to functions belonging to $K_{P}$ with degree $\lambda+2$. This is a contradiction, implying that if $i$ is odd, $c_{i}$ must be either 1 or $\lambda$. Analogously, we have $c_{i}$ must be either 1 or $\lambda$ if $i$ is even. Hence, we conclude that $c_{i}\in\{1,\lambda\}$ for all $i=1,\ldots,n$.
Suppose that $c_{1}=1$. Then, in view of Lemma 7.4, the functions $f$ and $g$ on the subset$\{X_{0},X_{1}\}$ defined by $f(X_{0})=1,f(X_{1})=\lambda+1$ and $g(X_{0})=2,g(X_{1})=\lambda+1$ can be extended two distinct functions in $K_{P}$ with a degree of $\lambda+2$. This is a contradiction, thereby proving that $c_{1}=\lambda$. Similarly, we can show that $c_{n}=1$.
It remains to show that there is no $22\leq i\leq n-$ 2 such that $c_{i}=c_{i+1}=1$. Suppose on the contrary that $c_{i}=c_{i+1}=1$ for some $22\leq i\leq n-$ 2. If $i$ is odd, according to Lemma 7.4, there are functions $f$ and $g$ of degree $\lambda+2$ such that $f(X_{b_{i-1}})=1,f(X_{b_{i}})=2$ and $f(X_{b_{i+1}})=1$, and that $g(X_{b_{i-1}})=1,g(X_{b_{i}})=\lambda+1$ and $g(X_{b_{i+1}})=1$. This is impossible. Similarly, the case when $i$ is even is also impossible.
Proposition 7.7.
Let $P$ be a basic poset of type $(c_{1},\ldots,c_{n})$. Then the following statements are equivalent:
$\mathrm{(1)}$ the chain algebra $\mathbb{K}[C_{P}]$ is nearly Gorenstein.
$\mathrm{(2)}$ the chain algebra $\mathbb{K}[C_{P}]$ is Gorenstein.
$\mathrm{(3)}$ it holds that $c_{1}=\cdots=c_{n}=\lambda.$
Proof.
$(1)\Leftrightarrow(2)$ By Theorem 4.2, the chain polytope $D_{P}$ of $P$ is indecomposable, and by Proposition 4.1, $\mathbb{K}[C_{P}]$ coincides with the Ehrhart ring of $D_{P}$. Now, the result follows from [6, Theorem 32].
(2) $\Rightarrow$ (3) Suppose on the contrary that $c_{1}=\cdots=c_{n}=\lambda$ does not hold. By lemma 7.6, there is $1<k<n$ such that $c_{k}=1$ and $c_{k-1}=c_{k+1}=\lambda\geq 2$. We may assume $k$ is even without loss of generality. By Lemma 7.4, there is at least one function $\gamma\in K_{P}$ of degree $12\lambda+$ 1 satisfying:
$\gamma(X_{b_{k-2}})=1,\gamma(X_{b_{k-2}+1})=2,\ldots,\gamma(X_{b_{k-1}})=\lambda+1,\gamma(X_{b_{k}})=\lambda,$
and
$\gamma(X_{b_{k}+1})=\lambda+1,\gamma(X_{b_{k}+1})=\lambda+2,\ldots,\gamma(X_{b_{k+1}})=2\lambda.$
On the other hand, since $\mathbb{K}[C_{P}]$ is Gorenstein, there is an unique function in $K_{P}$, say $f$, with degree $\lambda+2$ such that $K_{P}=f+C_{P}$. Note that $f(X_{b_{k-1}})=\lambda+1,f(X_{b_{k}})=1$, it follows that $(\gamma-f)(X_{b_{k-1}})=0$ and $(\gamma-f)(X_{b_{k}})=\lambda-1$. In view of Lemma 7.3, we have $\gamma-f$ does not belong to $C_{P}$ , a contradiction.
(3) $\Rightarrow$ (2) Let $f$ be the unique function in $K_{P}$ with $\deg(f)=\lambda+2$. Then the sequence $f(X_{0}),f(X_{1}),\ldots,f(X_{r})$ is as follows:
$1,2,\ldots,\lambda,\lambda+1,\lambda,\lambda-1,\ldots,1,2\ldots.$
From this sequence, we observe that $f(X_{i})-f(X_{i+1})\in\{1,-1\}$ for all $i=0,\ldots,r-1$. To prove $\mathbb{K}[C_{P}]$ is Gorenstein, it suffices to show that for any $g\in K_{P}$, the function $g-f$ belongs to $C_{P}$. Take an arbitrary function $g$ from $K_{P}$. According to Lemma 7.3, we have for all $i=0,\ldots,r-1$, $f(X_{i})$ and $g(X_{i})$ have the same ordering. That is, $f(X_{i})>f(X_{i+1})$ if and only if $g(X_{i})>g(X_{i+1})$, and vice versa. Since $g(X_{i})>0$ for all $i=0,\ldots,r$, it follows that $g(X_{i})\geq f(X_{i})$ for all $i$, and so $h:=g-f$ is a function from $P$ to $\mathbb{N}_{0}$.
Now, if $g(X_{i+1})>g(X_{i})$, then $g(X_{i+1})-g(X_{i})\geq f(X_{i+1})-f(X_{i})=1$, and so $h(X_{i+1})\geq h(X_{i})$. Similarly, if $g(X_{i+1})<g(X_{i})$, then $g(X_{i})-g(X_{i+1})\geq f(X_{i})-f(X_{i+1})=1$, and so $h(X_{i+1})\leq h(X_{i})$. Therefore, $h$ and $g$ share the same monotonicity, and so $h$ belongs to $C_{P}$ by Lemma 7.3, as desired.
Recalling [11, Theorem 2.7], we have if $R$ is the Segre product $R_{1}\#R_{2}\#\cdots\#R_{n}$ of Gorenstein standard graded $\mathbb{K}$-algebras $R_{i}$ with $\mathbf{a}(R_{i})=-a_{i}<0$ for $i=1,\ldots,n$, then $R$ is Gorenstein if and only if $a_{1}=\cdots=a_{n}$. Furthermore, $R$ is nearly Gorenstein if and only if $|a_{i}-a_{j}|\leq 1$ for all $1\leq i,j\leq n$.
Example 7.8.
Let $P$ be a pure poset of rank $r$, where $G_{i}(P)$ is a complete bipartite graph for all $i=0,\ldots,r-1$. Then $P$ is the ordinal sum of the subposets $P_{0},P_{1},\ldots,$ and $P_{r}$. Since $P_{i}$ is an anti-chain for each $i=0,\ldots,r$, $\mathbb{K}[C_{P_{i}}]$ is a polynomial ring with $|P_{i}|$ variables. In particular, $\mathbb{K}[C_{P_{i}}]$ is Gorenstein with $a(\mathbb{K}[C_{P_{i}}])=-|P_{i}|$. Therefore, according to [11, Theorem 2.7], we have $C_{P}$ is Gorenstein if and only if $|P_{0}|=\cdots=|P_{r}|$. Additionally, $C_{P}$ is nearly Gorenstein if and only if there is an integer $d\geq 1$ such that $|P_{i}|$ is equal to either $d$ or $d+1$ for each $i=0,\ldots,r$.
We are now ready to present the last main result of the paper.
Theorem 7.9.
Let $P$ be a pure poset with a width of 2, and denote by $\overline{P}$ the poset derived from $P$ by eliminating the disconnected bipartite graphs $G_{i}(P)$. Suppose that $\overline{P}$ can be expressed as the ordinal sum of basic posets $P_{1},\ldots,P_{s}$ and posets $Q_{1},\ldots,Q_{t}$ with rank 0 and width 2. Furthermore, assume that $P_{i}$ for $i=1,\ldots,s$ belongs to type $(c_{i,1},\ldots,c_{i,n_{i}})$.
$\mathrm{(1)}$ If $s=0$, then $\mathbb{K}[C_{P}]$ is always Gorenstein.
$\mathrm{(2)}$ If $t>0$ and $s>0$, then $\mathbb{K}[C_{P}]$ is not Gorenstein. Moreover, $\mathbb{K}[P]$ is nearly Gorenstein if and only if $c_{i,j}=1$ for all $i=1,\ldots,s,j=1,\ldots,n_{i}$.
$\mathrm{(3)}$ If $t=0$, then the following statements hold:
$\mathrm{(a)}$ $\mathbb{K}[C_{P}]$ is Gorenstein if and only if there is a positive integer $\lambda$ such that $c_{i,j}=\lambda$ for all $i=1,\ldots,s,j=1,\ldots,n_{i}$.
$\mathrm{(b)}$ $\mathbb{K}[C_{P}]$ is nearly Gorenstein if and only if there is a positive integer $\lambda$ and a subset $A\subseteq[1,s]$ such that $c_{i,j}=\lambda$ for $i\in A$ and $j=1,\ldots,n_{i}$, while $c_{i,j}=\lambda+1$ for $i\in[1,s]\setminus A$ and $j=1,\ldots,n_{i}$.
Proof.
Note that for each $i$, $\mathbb{K}[C_{Q_{i}}]$ is a Gorenstein algebra with a-invariant -2. Therefore, these conclusions are drawn by combining Propositions 7.7, and [11, Theorem 2.7].
We remark that conclusion (1) of the above theorem is actually a special case of Example 7.8.
Acknowledgment: This project is supported by NSFC (No. 11971338). We would like to express our sincere gratitude to Dr. Sora Miyashita, whose comments on the initial version of this paper brought us to the reference [6], resulting in the introduction and the exploration of the chain polytope in this revised version.